### v1

# Temporal Overlapping Prediction: A Self-supervised Pre-training Method for LiDAR Moving Object Segmentation

Abstract
Moving object segmentation (MOS) on LiDAR point clouds is crucial for autonomous systems like self-driving vehicles. Previous supervised approaches rely heavily on costly manual annotations, while LiDAR sequences naturally capture temporal motion cues that can be leveraged for self-supervised learning. In this paper, we propose Temporal Overlapping Prediction (TOP), a self-supervised pre-training method that alleviate the labeling burden for MOS. TOP explores the temporal overlapping points that commonly observed by current and adjacent scans, and learns spatiotemporal representations by predicting the occupancy states of temporal overlapping points. Moreover, we utilize current occupancy reconstruction as an auxiliary pre-training objective, which enhances the current structural awareness of the model. We conduct extensive experiments and observe that the conventional metric Intersection-over-Union (IoU) shows strong bias to objects with more scanned points, which might neglect small or distant objects. To compensate for this bias, we introduce an additional metric called $\text{mIoU}_{\text{obj}}$ to evaluate object-level performance. Experiments on nuScenes and SemanticKITTI show that TOP outperforms both supervised training-from-scratch baseline and other self-supervised pre-training baselines by up to 28.77% relative improvement, demonstrating strong transferability across LiDAR setups and generalization to other tasks. Code and pre-trained models will be publicly available upon publication.
1 Introduction
Moving Object Segmentation (MOS) on LiDAR data aims to segment moving points from the entire LiDAR scan and provides autonomous systems [9, 33, 34, 50, 54] with important clues in localization and dynamic obstacle avoidance, which enables better planning of collision-free trajectories in dynamic scenes. However, per-point manual labeling of moving objects [9, 3, 6] is both costly and labor-intensive. Meanwhile, self-supervised pre-training achieves significant success in Natural Language Processing (NLP) [13, 30, 19, 40, 5], Computer Vision (CV) [51, 17, 18, 14, 41], and 3D Vision [59, 39, 55, 58], enabling the learning of transferable embeddings from unlabeled raw data. This motivates us to pre-train a backbone using raw LiDAR data in a self-supervised manner and then fine-tune it for MOS, which has limited expensive labeled data.
The MOS task heavily relies on temporal information, including the positions of an object at different times and the changing relative positions of an object to its surroundings. In the literature involving temporal information of LiDAR data, an intuitive approach is to utilize LiDAR observation forecasting as a self-supervised pre-training method for MOS. The forecasting task includes point cloud forecasting  [53, 52, 32, 60, 23] and occupancy forecasting [1], which focus on predicting future scans or occupancy maps based on past and current LiDAR scans and poses. However, utilizing forecasting as a pre-training method for MOS presents certain limitations: (1) The forecasting task is more complex and advanced than MOS. Forecasting a static object mainly requires geometric information to reconstruct the unseen parts of the object from a novel sensor view. For moving objects, beyond geometry, it is essential to extract temporal features such as the object’s kinematics and interactions between objects and their surroundings. Hence, relying on forecasting as a pre-training step for MOS could be overly demanding. (2) Existing methods [23, 1] inevitably learn a “flow” when using raw future observations as supervision. This complex “flow” information represents the relationship between past observations’ embeddings and future observations. Current approaches either implicitly learn this flow through simple feature broadcasting [23] or explicitly predict it [1], lacking direct supervision similar to the scene flow task [61, 62, 29]. Thus, the learned “flow” tends to be noisy. Relying on this noisy “flow” for deterministic future observation predictions (as opposed to the generative approach [60]) can lead to network overfitting, thereby reducing its transferability to target downstream tasks.
Figure 1: Illustration of temporal overlapping points in a realistic scenario. The color of points represents the occupancy state, and the index marked on the points indicates the observed time. For the two temporal overlapping points in the past and one in the future, their temporal occupancy states differ from the current occupancy states.
To overcome these two limitations and incorporate temporal information for MOS self-supervised pre-training, we propose Temporal Overlapping Prediction. TOP explores the temporal occupancy states of temporal overlapping points observed by both the current scan and adjacent past or future scans for pre-training. The rationale for this lies in the potential for moving objects to cause the current occupancy states of temporal overlapping points to differ from the states at other observed times, as illustrated in Figure 1. We first pre-process the spatial positions and temporal occupancy states of the overlapping points with beam divergence [31], which is a physical characteristic of LiDAR. This LiDAR nature helps avoid discretization errors caused by manually setting a fixed voxel resolution. Then, we pre-train the encoder by predicting the pre-processed temporal occupancy states, which simplifies the task compared to directly forecasting LiDAR observations. Based on a distinct temporal correlation between LiDAR beams at different times established by the pre-processing, our pre-training factors out either implicit or explicit learning of a noisy “flow”. Additionally, since understanding current states is also important for MOS, we add a reconstruction term for optimization.
We assess the MOS performance on the nuScenes [6] and SemanticKITTI [3] datasets. Previous evaluations of nuScenes MOS in [34, 46] directly convert object attributes to MOS labels. However, these attributes do not precisely reflect the motion state of the object, and not all objects have attributes. Thus, we combine object attributes with object speed calculated based on bounding boxes to obtain refined nuScenes MOS labels. In the experiments, we observe that the conventional metric Intersection-over-Union (IoU) for MOS exhibits a significant bias towards objects that have a larger number of scanned points, potentially overlooking smaller or more distant objects. Thus, we propose an additional object-level evaluation metric $\text{mIoU}_{\text{obj}}$ to take these objects into account.
We conducted few-shot fine-tuning experiments on the nuScenes dataset with different amounts of labeled data. Results show that TOP achieves improvements over the supervised training-from-scratch baseline [33] and outperforms other self-supervised pre-training baselines, such as state-of-the-art (SOTA) LiDAR observation forecasting methods [23, 1] and the occupancy-based 3D pre-training method [4]. The results of our few-shot cross-dataset experiments on the SemanticKITTI dataset further highlight the superior transferability of TOP pre-training compared to various baselines. Additionally, we fine-tune the pre-trained backbone on the nuScenes semantic segmentation task in a few-shot setting, demonstrating the potential generalization capability of TOP across more tasks.
2 Related Works
Moving Object Segmentation. Existing supervised MOS methods can be grouped into projection-based and voxel-based approaches. The early MOS methods [9, 24, 45, 11] project 3D LiDAR data to range images and embed the range residual images by 2D backbone. These methods suffer from imprecise boundaries when re-projecting 2D outputs to 3D point clouds [34]. Recent methods [46, 33, 34, 50] utilize the sparse convolution backbone to directly embed voxelized point clouds. 4DMOS [33] first employs sparse 4D UNet [12] and makes it the SOTA backbone for MOS. For unsupervised approaches, several scene flow methods [48, 2, 16, 44] decouple the per-point flow to an ego-motion flow for static points and a non-rigid flow for moving points, which provides a byproduct of MOS [33]. However, these approaches using only two adjacent scans, encounter challenges in identifying slowly moving objects [33]. MOTS [25] proposes a label-free method that only works in stationary scenes. 4dNDF [64] encodes the 4D scene into an implicit neural map, which can be used to distinguish the static map and dynamic objects. However, its application in real-time is limited. In addition, M-detector [54] proposes a network-free approach based on the occlusion principle. It requires manual tuning of numerous hyper-parameters when using different LiDARs or deploying in varying scenes. Moreover, to reduce the cost of manual labeling, several methods [10, 28] propose a pipeline that first generates coarse labels automatically and then refines them manually. In contrast to previous works on MOS, we employ self-supervised pre-training to enhance the few-shot MOS performance.
3D Self-supervised Pre-training. Effective self-supervised pre-training methods in NLP [13, 40, 5] and CV [18] have been extended to 3D data. Some methods [59, 15, 7] leverage BERT-like [13] or GPT-like [40, 5] pre-training on object CAD point clouds. Contrastive learning methods [55, 42, 38, 22, 27, 8] employ various transformations or augmentations to construct point/instance/scene-level contrastive loss, learning geometric or semantic information without manual labels. Additionally, Masked Auto-Encoder (MAE)-based methods [39, 63, 20, 57, 47, 36, 26, 56] extend MAE [18] to 3D point clouds. These methods train an encoder by predicting a certain ratio of masked point cloud patches, which transfers well on object detection and semantic segmentation tasks. Subsequent works [21, 65, 58] apply NeRF-like [35] volumetric neural rendering to leverage multi-modal inputs and bridge the gap between 2D and 3D data. In summary, existing 3D self-supervised pre-training works lack temporal information and thus are unsuitable for MOS.
LiDAR Observation Forecasting. Existing works that involve temporal information of LiDAR data mainly focus on LiDAR observation forecasting, including point cloud and occupancy forecasting. Previous works on point cloud forecasting [53, 52, 32] directly predict the raw future LiDAR scans based on past ones. A recent study, 4Docc [23], claims the potential of forecasting as a scalable self-supervised task. Instead of directly predicting raw point clouds, 4Docc renders the future point clouds on a predicted occupancy map using known sensor intrinsics and extrinsics, enabling the network to focus on learning the world rather than sensor characteristics. Another generative approach, Copilot4D [60] introduces the concept of the world model. It tokenizes LiDAR scans and predicts future scans through discrete diffusion. However, the learned world knowledge is difficult to apply to downstream tasks due to the nature of heavy-decoder. ALSO [4] is a single-scan method that reconstructs the occupancy of a local surface and inspires UnO [1]. UnO is the most recent occupancy forecasting method that predicts the binary occupancy class of points sampled along future LiDAR rays. In this paper, we pre-train the backbone using the proposed TOP method instead of future observation forecasting.
3 Methodology
Figure 2: The overall pipeline. Pre-processing: (a) Coplanarity condition of two beams. (b) If the spatial angle between two beams $\alpha_{i,j}$ exceeds the beam divergence angle $\theta_{\text{dvg}}$, we sample the beams’ intersection point as the temporal overlapping point. The red and blue areas indicate beam divergence. (c) When $\alpha_{i,j}$ is less than $\theta_{\text{dvg}}$, the intersection segment is calculated to sample temporal overlapping points within it. Pre-training: (d) We use a sparse 4D UNet to embed the input scans. For the temporal overlapping prediction objective $\mathcal{L}_{\text{overlap}}$, the occupancy state of a temporal overlapping point is predicted based on its positional encoding and the embedding of the current beam hit-point. Similarly, the occupancy state of the currently sampled point is also predicted for the current reconstruction objective $\mathcal{L}_{\text{recon}}$.
In this section, we introduce the self-supervised pre-training method TOP. Our method aims to pre-train an encoder for the downstream MOS task. The pre-training goal is to predict the occupancy states of temporal overlapping points and to reconstruct the current occupancy. We first introduce notations and problem formulation in Sec. 3.1. Then, we discuss the pre-processing of the temporal overlapping points in Sec. 3.2. Finally, we describe the pre-training objectives in Sec. 3.3.
3.1 Problem Formulation
Notations. To begin with, we define a LiDAR scan as $\mathbf{P}\in\mathbb{R}^{N\times 3}$, where $N$ is the number of points. We use subscript $i$ to indicate the $i^{th}$ point in $\mathbf{P}$, where each point $\mathbf{p}_{i}\in\mathbb{R}^{3}$ ($i=1,2,...,N$) contains $xyz$ location. $\mathbf{p}_{i}$ can be further described as a laser beam emitted from the sensor origin $\mathbf{a}\in\mathbb{R}^{3}$ with its normalized beam direction $\mathbf{d}_{i}\in\mathbb{R}^{3}$ and range $r_{i}$, leading to $\mathbf{p}_{i}=\mathbf{a}+r_{i}\mathbf{d}_{i}$. Although the laser beam is commonly modeled as an ideal ray, the real LiDAR beam exhibits a diverging radius as it travels, and this is referred to as beam divergence [31]. The beam divergence angle, denoted as $\theta_{\text{dvg}}$, is a constant for each type of LiDAR. When any part of the diverging beam hits an object, the LiDAR calculates the distance by time of flight and considers the beam center point as the hit-point, as shown in Fig. 3. To incorporate the temporal LiDAR sequence, we indicate scans at different timestamps with superscripts. For example, the scan at the current timestamp $t_{0}$ is $\mathbf{P}^{t_{0}}\in\mathbb{R}^{N\times 3}$. For adjacent past/future timestamps $\mathcal{T}=\{t_{-n}\ldots t_{-1},t_{1}\ldots t_{n}\}$, we have $\mathcal{P}_{\text{adj}}=\{\mathbf{P}^{t}\in\mathbb{R}^{N_{t}\times 3}\}_{t\in\mathcal{T}}$. The $j^{th}$ beam at time $t$, denoted as $\mathbf{d}_{j}^{t}$, can also be described by $\mathbf{p}_{j}^{t}=\mathbf{a}^{t}+r_{j}^{t}\mathbf{d}_{j}^{t}$.
Point Clouds Encoding. All scans in $\mathcal{P}_{\text{adj}}$ are transformed to the local coordinate system of $\mathbf{P}^{t_{0}}$ with the poses provided in the datasets. We take $n$ current and past scans as network inputs $\mathcal{P}_{\text{in}}=\{\mathbf{P}^{t_{-n+1}}\ldots\mathbf{P}^{t_{-1}},\mathbf{P}^{t_{0}}\}$. Each point in $\mathcal{P}_{\text{in}}$ is concatenated with its timestamp as a 4D point before being fed into the network. We apply a sparse 4D UNet encoder $g(\cdot)$ used in [12, 33, 34, 50] to embed the spatiotemporal information of the 4D inputs:
$\mathbf{F}=g(\mathcal{P}_{\text{in}}),$ (1)
where $\mathbf{F}\in\mathbb{R}^{N\times d}$ is the fully sparse features of the current scan $\mathbf{P}^{t_{0}}$ with a feature dimension of $d$.
Pre-training Loss. We incorporate temporal overlapping points and their temporal occupancy states to pre-train the encoder $g(\cdot)$. As described in Fig. 1, temporal overlapping points are located on current LiDAR beams in $\mathbf{P}^{t_{0}}$ and are also observed by adjacent past/future beams in $\mathcal{P}_{\text{adj}}$. Firstly, we denote the temporal overlapping points as $\mathbf{O}\in\mathbb{R}^{M\times 4}$, where $M$ is the number of points and each point contains $xyz$ position and the temporally observed timestamp $t$. Then, we use the LiDAR occupancy measurement model (introduced in Sec. 3.2) to describe the occupancy state labels of $\mathbf{O}$, denoted as $\mathbf{S}\in\mathbb{R}^{M\times 3}$. Each label is a three-dimensional one-hot vector indicating the occupancy states “free”, “occupied”, and “unknown”. The overall pre-training goal consists of temporal occupancy prediction and current occupancy reconstruction, which is described as:
$\mathcal{L}=\mathcal{L}_{\text{overlap}}(\mathbf{O},\mathbf{S},\mathbf{F})+\mathcal{L}_{\text{recon}}(\mathbf{P}^{t_{0}},\mathbf{F}).$ (2)
Figure 3: LiDAR beam divergence model: The bottom-left blue cone illustrates a real diverging beam, where the beam radius expands as the traversed distance increases. When any point within the beam spot hits an obstacle (indicated by the real hit-point colored in red), the center point of the beam spot (colored in black) is considered as the measured point. LiDAR occupancy measurement model: The occupancy state is confidently classified as free before reaching the beam hit-point. Then, the confidence level exponentially decays from the beam hit-point onward. A small area of high confidence after the beam hit-point is categorized as the occupied class, while the remaining areas are deemed unknown.
3.2 Temporal Overlapping Pre-processing
As computing temporal overlapping during pre-training can be inefficient, we pre-process the LiDAR scans $\mathbf{P}^{t_{0}}$ and $\mathcal{P}_{\text{adj}}$ to produce temporal overlapping points $\mathbf{O}$ and their occupancy states $\mathbf{S}$. We first identify coplanar beam pairs from $\mathbf{P}^{t_{0}}$ and $\mathcal{P}_{\text{adj}}$, as coplanarity is a necessary condition for beams to have overlapping observations, as illustrated in Fig. 2(a). Then, we compute the intersection of coplanar beam pairs and sample temporal overlapping points. Depending on the spatial angle between two beams, we split the computation and sampling into two scenarios, shown in Fig. 2(b) and Fig. 2(c), respectively. We take an example beam pair for a clearer explanation: the $i^{th}$ beam at time $t_{0}$ and the $j^{th}$ beam at time $t$ ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$), emitted from ($\mathbf{a}^{t_{0}}$, $\mathbf{a}^{t}$) and hitting ($\mathbf{p}_{i}^{t_{0}}$, $\mathbf{p}_{j}^{t}$). When a beam pair ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$) generates the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$, we record the $(i,j,k)$ correspondence for subsequent usage in Sec. 3.3.
Coplanarity. We define a reference plane formed by ($\mathbf{a}^{t_{0}}$, $\mathbf{a}^{t}$, $\mathbf{p}_{i}^{t_{0}}$), illustrated as the yellow triangle in Fig. 2(a). The unit normal vector of this plane is:
$\mathbf{n}_{i}^{t}=\mathbf{d}_{i}^{t_{0}}\times\frac{\mathbf{a}^{t}}{||\mathbf{a}^{t}||}.$ (3)
The space angle between beam $\mathbf{d}_{j}^{t}$ and the plane is:
$\mathbf{\theta}_{i,j}=\arccos{(\mathbf{n}_{i}^{t}\cdot\mathbf{d}_{j}^{t})}-\frac{\pi}{2}.$ (4)
Based on the physical beam divergence, if $|\mathbf{\theta}_{i,j}|\leq\frac{1}{2}\theta_{\text{dvg}}$, we assume the two beams ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$) are coplanar. All coplanar beam pairs are split into two scenarios based on the value of their spatial angle $\alpha_{i,j}=\arccos{(\mathbf{d}_{i}^{t_{0}}\cdot\mathbf{d}_{j}^{t})}$.
Scenario-1 ($\alpha_{i,j}>\theta_{\text{dvg}}$). This scenario describes most overlapping cases where the spatial angle between two beams exceeds the beam divergence angle, as shown in Fig. 2(b). The beam intersection is small, thus being considered as a point $\mathbf{q}_{i,j}$, which is the intersection of the two beams’ centerlines. The $\mathbf{q}_{i,j}$ is solved from the parameter equations of the two beams:
$\mathbf{q}_{i,j}=\frac{\mathbf{a}^{t}\times\mathbf{d}_{j}^{t}\cdot\mathbf{m}_{i,j}}{\mathbf{m}_{i,j}\cdot\mathbf{m}_{i,j}}\cdot\mathbf{d}_{i}^{t_{0}},$ (5)
where $\mathbf{m}_{i,j}=\mathbf{d}_{i}^{t_{0}}\times\mathbf{d}_{j}^{t}$ is the unit common perpendicular line of the two beams. In this scenario, the intersection point $\mathbf{q}_{i,j}$ is selected as the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$.
Scenario-2 ($\alpha_{i,j}\leq\theta_{\text{dvg}}$). When the spatial angle between two beams is less than the beam divergence, the intersection of the two beams can no longer be considered as a point but an area, as shown in Fig. 2(c). In this scenario, we compute the intersection segments (purple dashed lines) on the centerlines for subsequent sampling. Firstly, we obtain the endpoint $\mathbf{q}_{i,j}$ from the beam intersection similar to Scenario-1. Then, we approximate the starting point $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ as the projection from a point $\mathbf{b}$ to the two beams. Here, $\mathbf{b}$ is the crossed point of the boundaries of the two beams. Given that the spatial angle $\alpha_{i,j}$ is small, we make two approximations to compute the starting points: (1) We approximate that the two intersection segments share the same length. (2) We approximate that the distance between the two starting points $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ is equal to the sum of the beam radius at the two points. The detailed formulas for computing $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ are in the appendix.
Since the overlap between two beams is an area, we sample 5 temporal overlapping points inside the intersection segment, as shown in Fig. 2(c) right: $\mathbf{o}_{k_{1}}$, the beam hit point $\mathbf{p}_{i}^{t_{0}}$; $\mathbf{o}_{k_{2}}$, the projection of $\mathbf{p}_{j}^{t}$ on $\mathbf{d}_{i}^{t_{0}}$; $\mathbf{o}_{k_{3}}$, the center of ($\mathbf{o}_{k_{1}}$, $\mathbf{o}_{k_{2}}$); $\mathbf{o}_{k_{4}}$, the center of ($\mathbf{o}_{k_{1}}$, $\mathbf{q}_{i,j}$); $\mathbf{o}_{k_{5}}$, the center of ($\mathbf{o}_{k_{2}}$, $\mathbf{q}_{i,j}$). Their occupancy states at timestamp $t$ are given by the states of their projection points on $\mathbf{d}_{j}^{t}$.
LiDAR Occupancy Measurement Model. The LiDAR occupancy measurement model [1, 43] partitions the space between the sensor origin to infinity along the beam direction into three areas with different occupancy states, as shown in Fig. 3. For the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$ generated from the overlap beams pair ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$), the confidence level $w_{\text{conf}}$ of the measured occupancy state of $\mathbf{o}_{k}$ is 100% when $\mathbf{o}_{k}$ lies inside the physically traversed areas of the beam (blue area in Fig. 3). It then exponentially decays after the hit-point $\mathbf{p}$, as described below:
$w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t})=\begin{cases}1,\ \text{if}\ 0\leq||\mathbf{o}_{k}||\leq||\mathbf{p}_{j}^{t}||,\\ e^{{-(||\mathbf{o}_{k}||-||\mathbf{p}_{j}^{t}||)}},\ \text{if}\ ||\mathbf{o}_{k}||>||\mathbf{p}_{j}^{t}||.\\ \end{cases}$ (6)
The occupancy state of $\mathbf{o}_{k}$ is denoted as $\mathbf{s}_{k}$, which is described as:
$\mathbf{s}_{k}=\begin{cases}[1,0,0],\ \text{if}\ 0\leq||\mathbf{o}_{k}||<||\mathbf{p}_{j}^{t}||,\\ [0,1,0],\ \text{if}\ ||\mathbf{o}_{k}||=||\mathbf{p}_{j}^{t}||\ \text{or}\ \lambda_{\text{occ}}\leq w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t}),\\ [0,0,1],\ \text{if}\ w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t})<\lambda_{\text{occ}},\\ \end{cases}$ (7)
where $\mathbf{s}_{k}$ is a one-hot vector indicating “free”, “occupied” and “unknown” occupancy states, and $\lambda_{\text{occ}}$ is the confidence level threshold of the occupied state.
3.3 Training Objectives
For self-supervised pre-training, the primary objective is temporal overlapping prediction, which involves predicting $\mathbf{s}_{k}$ from the $i^{th}$ current point’s feature $\mathbf{f}_{i}$ and the positional information of $\mathbf{o}_{k}$ using a shallow MLP decoder $h(\cdot)$:
$\mathbf{u}_{k}=h(f(\mathbf{o}_{k})+\mathbf{f}_{i})\in\mathbb{R}^{3},$ (8)
where $\mathbf{u}_{k}$ is the predicted probabilities of three occupancy states, and $f(\mathbf{o}_{k})\in\mathbb{R}^{d}$ denotes the sine-cosine 4D positional encoding vector as described in [49]. Temporal overlapping prediction is described as a Cross-Entropy loss:
$\mathcal{L}_{\text{overlap}}=-\frac{1}{M}\sum\limits_{k=1}^{M}w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t})\cdot\mathbf{w}_{s}\cdot\mathbf{s}_{k}\cdot\log{(\mathbf{u}_{k})},$ (9)
where $\mathbf{w}_{s}$ is a constant weight vector for each occupancy state, and the $(i,j,k)$ correspondence is introduced in Sec. 3.2.
Additionally, we reconstruct the current occupancy using the input scans, which serves as a regularization term in the total loss. We randomly select $M^{\prime}$ occupied and free points on each current beam. For the $k^{th}$ point $\mathbf{c}_{k}$ sampled on the $i^{th}$ current beam, it includes the spatial position and timestamp $t_{0}$. The ground truth occupancy label is $\mathbf{s}_{k}$, and the predicted probability is $\mathbf{v}_{k}$. The reconstruction objective is also specified as a Cross-Entropy loss:
$\mathbf{v}_{k}$ $=h(f(\mathbf{c}_{k})+\mathbf{f}_{i})\in\mathbb{R}^{3},$ (10)
$\mathcal{L}_{\text{recon}}$ $=-\frac{1}{NM^{\prime}}\sum\limits_{k=1}^{NM^{\prime}}\mathbf{w}_{s}\cdot\mathbf{s}_{k}\cdot\log{(\mathbf{v}_{k})}.$ (11)
4 Evaluation Metrics
In this section, we propose a mean object-level metric denoted as $\text{mIoU}_{\text{obj}}$ to tackle the biased weighting issue in the conventional IoU metric. Previous methods [9, 33, 34, 50] have relied solely on IoU as the evaluation metric. However, IoU implicitly gives larger weight to objects with more scanned points due to its point-wise nature. Objects of varying geometric sizes and distances to the LiDAR may result in different numbers of scanned points. Nevertheless, these objects are equally important in practical MOS applications and should not be weighted based on the number of scanned points. To further illustrate the impact of this bias on MOS performance evaluation, we conduct a statistical analysis: we sort all moving objects in the nuScenes train-val split based on the number of scanned points, and then calculate the cumulative distribution of moving objects and scanned points. The results indicate that the top 75.22% of objects with the fewest scanned points only contribute to 15.49% of the total points, while 90.06% of objects contribute to 37.02% of the points, revealing that a few objects with more points heavily influence the conventional IoU metric. The cumulative distribution is provided in the appendix.
The proposed metric $\text{mIoU}_{\text{obj}}$ calculates the IoU for each moving object and computes the average, ensuring equal weighting for each moving object:
$\text{mIoU}_{\text{obj}}=\frac{1}{m}\sum\limits_{i=1}^{m}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}},$ (12)
where $m$ denotes the number of moving objects, $\text{TP}_{i}$ and $\text{FP}_{i}$ indicate the true positive and false positive classifications of the points belonging to $i^{th}$ moving object. False negatives are not involved in the $\text{mIoU}_{\text{obj}}$ as all points are classified as the moving (positive) class.
Additionally, in the nuScenes train-val split, ego vehicles’ moving points account for 76.72% of all moving points, which can heavily influence the IoU. However, in autonomous driving scenarios, we already know the movement of the ego vehicle and do not need to predict it. Therefore, we utilize a refined IoU metric called $\text{IoU}_{\text{w/o}}$, calculated as described in [9, 33, 34] but excluding the scanned ego vehicle points.
For a more precise and robust evaluation of the MOS performance, we use both metrics: $\text{mIoU}_{\text{obj}}$ assesses performance on all moving objects, while $\text{IoU}_{\text{w/o}}$ reflects overall performance, incorporating false negatives.
5 Experiments
5.1 Experiments Setup
Datasets. We pretrain the encoder on the nuScenes dataset [6], which utilizes a 32-beam Velodyne LiDAR for data collection. The beam divergence angle $\theta_{\text{dvg}}$ specified by the manufacturer is 0.003 rad. We perform few-shot experiments on the nuScenes dataset. Previous methods [34, 46] convert nuScenes object attributes into MOS labels. However, these attributes do not precisely describe the object’s motion state. Moreover, some objects lack attributes and will be overlooked during training or evaluation. To overcome this limitation, we refine MOS labels by incorporating object attributes and speed. Speed is computed from object bounding boxes and is used to classify motion states by static and moving speed thresholds. We introduce a speed range between these thresholds to categorize objects with unclear motion states into an unknown class, such as cars that are just starting or hesitant pedestrians. More details are available in the appendix. We also conduct few-shot cross-dataset transfer experiments on the SemanticKITTI [3] MOS benchmark [9], which employs a 64-beam Velodyne LiDAR for data collection.
Implementation Details. In our setup, we have $n=6$, where $\mathcal{P}_{\text{adj}}$ includes 6 past scans and 6 future scans, and the input sequence $\mathcal{P}_{\text{in}}$ comprises 6 scans. All scans are sampled at 2Hz. The confidence cutoff threshold for the occupied state is $\lambda_{\text{occ}}=0.9$. We apply data augmentation techniques from [37, 33] and discard points outside a bounded area based on [23, 1]. Our encoder is the sparse 4D UNet backbone, commonly used in MOS methods [33, 34]. The quantization size for the sparse backbone is 0.1m, and the feature dimension $d=128$. For the temporal overlapping prediction objective, all occupied class samples are included due to their scarcity, while 5 times the number of occupied samples are randomly selected for the free class and the same amount for the unknown class in each training iteration. As for the current occupancy reconstruction objective, we select 5 occupied points and 25 free points on each current beam ($M^{\prime}=30$). The class weight is $\mathbf{w}_{s}=[1,5,1]$ for free, occupied, and unknown class. We use a fixed random seed for all experiments to ensure reproducibility. Detailed optimizer settings are in the appendix.
Baselines. We use the SOTA supervised MOS method 4DMOS [33] as the training-from-scratch baseline. Additionally, we implement the SOTA LiDAR observation forecasting methods 4Docc [23] and UnO [1], along with the occupancy-based 3D method ALSO [4], as the self-supervised pre-training baselines. All baselines utilize the same sparse 4D UNet [12, 33] encoder with their respective original decoders.
5.2 Few-shot Results on nuScenes
We perform few-shot MOS experiments on the nuScenes dataset. All methods are first pre-trained on full sequences from the nuScenes train split. Then the pre-trained encoder of the last epoch is chosen and fine-tuned with labeled MOS data using binary Cross-Entropy loss. We randomly select 5%/10%/20%/50% of sequences from the nuScenes train split for MOS fine-tuning. To assess the convergence performance, all models are trained for a sufficient number of epochs until convergence. As it is not possible to assign weights to the two metrics to select a single “best” model, we can only choose two best models separately based on each metric. We present the test results on full sequences from the nuScenes val split in Tab. 1.
Our pre-training method improves training-from-scratch baseline across all fine-tuning data amounts. Our method achieves significant improvement in $\text{mIoU}_{\text{obj}}$, up to 12.17% relative improvement from 24.98% to 28.03% $\text{mIoU}_{\text{obj}}$ when using 10% data, while the improvement in $\text{IoU}_{\text{w/o}}$ is less pronounced. This is because we detect more moving objects with less number of points (such as pedestrians or cycles), which account for a small proportion of the total points. It is noteworthy that UnO also shows some increases on $\text{mIoU}_{\text{obj}}$, while the $\text{IoU}_{\text{w/o}}$ all decreases. This result indicates that the performance of pre-training by UnO is not consistent, and it introduces more false negatives when detecting moving objects. This will lead to erroneous moving object alerts in MOS applications and cause incorrect path planning. In contrast, our method demonstrates consistent results in most cases. We visualize some results in Fig. 4.
Data Pretrain Best $\textbf{mIoU}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
$\textbf{mIoU}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{mIoU}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
5% No 24.27 34.96 23.20 35.05
ALSO $\text{18.68}^{\text{{\color[rgb]{0.68,0,0}-5.59}}}$ $\text{31.65}^{\text{{\color[rgb]{0.68,0,0}-3.31}}}$ $\text{17.57}^{\text{{\color[rgb]{0.68,0,0}-5.63}}}$ $\text{31.97}^{\text{{\color[rgb]{0.68,0,0}-3.07}}}$
4Docc $\text{22.14}^{\text{{\color[rgb]{0.68,0,0}-2.13}}}$ $\text{33.39}^{\text{{\color[rgb]{0.68,0,0}-1.57}}}$ $\text{21.02}^{\text{{\color[rgb]{0.68,0,0}-2.17}}}$ $\text{33.81}^{\text{{\color[rgb]{0.68,0,0}-1.24}}}$
Uno $\text{25.23}^{\text{{\color[rgb]{0,0.88,0}+0.96}}}$ $\text{34.31}^{\text{{\color[rgb]{0.68,0,0}-0.65}}}$ $\text{22.62}^{\text{{\color[rgb]{0.68,0,0}-0.58}}}$ $\text{34.90}^{\text{{\color[rgb]{0.68,0,0}-0.15}}}$
Ours $\textbf{25.75}^{\textbf{{\color[rgb]{0,0.88,0}+1.48}}}$ $\textbf{35.94}^{\textbf{{\color[rgb]{0,0.88,0}+0.98}}}$ $\textbf{25.75}^{\textbf{{\color[rgb]{0,0.88,0}+2.56}}}$ $\textbf{35.94}^{\textbf{{\color[rgb]{0,0.88,0}+0.90}}}$
10% No 24.98 36.44 22.98 38.06
ALSO $\text{27.32}^{\text{{\color[rgb]{0,0.88,0}+2.34}}}$ $\text{35.74}^{\text{{\color[rgb]{0.68,0,0}-0.70}}}$ $\text{22.47}^{\text{{\color[rgb]{0.68,0,0}-0.51}}}$ $\textbf{38.33}^{\textbf{{\color[rgb]{0,0.88,0}+0.27}}}$
4Docc $\text{24.31}^{\text{{\color[rgb]{0.68,0,0}-0.67}}}$ $\text{36.01}^{\text{{\color[rgb]{0.68,0,0}-0.43}}}$ $\text{23.83}^{\text{{\color[rgb]{0.68,0,0}+0.86}}}$ $\text{37.46}^{\text{{\color[rgb]{0.68,0,0}-0.70}}}$
Uno $\text{26.38}^{\text{{\color[rgb]{0,0.88,0}+1.40}}}$ $\text{36.56}^{\text{{\color[rgb]{0,0.88,0}+0.12}}}$ $\text{22.00}^{\text{{\color[rgb]{0.68,0,0}-0.98}}}$ $\text{37.17}^{\text{{\color[rgb]{0.68,0,0}-0.89}}}$
Ours $\textbf{28.03}^{\textbf{{\color[rgb]{0,0.88,0}+3.04}}}$ $\textbf{36.95}^{\textbf{{\color[rgb]{0,0.88,0}+0.51}}}$ $\textbf{24.69}^{\textbf{{\color[rgb]{0,0.88,0}+1.72}}}$ $\text{38.06}^{\text{{\color[rgb]{0,0.88,0}+0.00}}}$
20% No 25.59 44.25 24.52 45.05
ALSO $\text{26.68}^{\text{{\color[rgb]{0,0.88,0}+1.08}}}$ $\text{40.84}^{\text{{\color[rgb]{0.68,0,0}-3.41}}}$ $\text{25.81}^{\text{{\color[rgb]{0,0.88,0}+1.28}}}$ $\text{42.33}^{\text{{\color[rgb]{0.68,0,0}-2.72}}}$
4Docc $\text{26.01}^{\text{{\color[rgb]{0,0.88,0}+0.42}}}$ $\text{42.21}^{\text{{\color[rgb]{0.68,0,0}-2.04}}}$ $\text{24.58}^{\text{{\color[rgb]{0,0.88,0}+0.06}}}$ $\text{43.11}^{\text{{\color[rgb]{0.68,0,0}-1.94}}}$
Uno $\text{28.14}^{\text{{\color[rgb]{0,0.88,0}+2.55}}}$ $\text{43.63}^{\text{{\color[rgb]{0.68,0,0}-0.62}}}$ $\text{26.78}^{\text{{\color[rgb]{0,0.88,0}+2.26}}}$ $\text{43.97}^{\text{{\color[rgb]{0.68,0,0}-1.08}}}$
Ours $\textbf{28.45}^{\textbf{{\color[rgb]{0,0.88,0}+2.86}}}$ $\textbf{44.30}^{\textbf{{\color[rgb]{0,0.88,0}+0.05}}}$ $\textbf{27.29}^{\textbf{{\color[rgb]{0,0.88,0}+2.76}}}$ $\text{44.52}^{\text{{\color[rgb]{0.68,0,0}-0.53}}}$
50% No 29.80 49.29 29.80 49.29
ALSO $\text{30.72}^{\text{{\color[rgb]{0,0.88,0}+0.91}}}$ $\text{47.34}^{\text{{\color[rgb]{0.68,0,0}-1.95}}}$ $\text{28.06}^{\text{{\color[rgb]{0.68,0,0}-1.75}}}$ $\text{48.10}^{\text{{\color[rgb]{0.68,0,0}-1.19}}}$
4Docc $\text{29.20}^{\text{{\color[rgb]{0.68,0,0}-0.60}}}$ $\text{46.73}^{\text{{\color[rgb]{0.68,0,0}-2.56}}}$ $\text{28.19}^{\text{{\color[rgb]{0.68,0,0}-1.62}}}$ $\text{48.19}^{\text{{\color[rgb]{0.68,0,0}-1.10}}}$
Uno $\text{29.54}^{\text{{\color[rgb]{0.68,0,0}-0.26}}}$ $\text{48.59}^{\text{{\color[rgb]{0.68,0,0}-0.70}}}$ $\text{28.81}^{\text{{\color[rgb]{0.68,0,0}-1.00}}}$ $\text{48.76}^{\text{{\color[rgb]{0.68,0,0}-0.53}}}$
Ours $\textbf{30.58}^{\textbf{{\color[rgb]{0,0.88,0}+0.78}}}$ $\textbf{49.80}^{\textbf{{\color[rgb]{0,0.88,0}+0.51}}}$ $\textbf{30.58}^{\textbf{{\color[rgb]{0,0.88,0}+0.78}}}$ $\textbf{49.80}^{\textbf{{\color[rgb]{0,0.88,0}+0.51}}}$
Table 1: Few-shot MOS results on the nuScenes dataset. The best performance in each column is in bold. The metric used for model selection is in bold in the table header.
5.3 Transferring Results on SemanticKITTI
To evaluate the transferability of the learned spatiotemporal representations, we fine-tune the pre-trained 4D UNet encoder on the SemanticKITTI dataset. We train the model on 10%/20%50% random samples from SemanticKITTI train set for enough epochs until convergence. As demonstrated in Tab. 2, TOP shows the best transferability on different LiDAR setups among all baselines. Fig. 5 displays the results on the SemanticKITTI dataset.
Data Pretrain Best $\textbf{mIoU}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
$\textbf{mIoU}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{mIoU}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
10% No 37.22 41.22 33.97 42.81
ALSO $\text{9.21}^{\text{{\color[rgb]{0.68,0,0}-28.01}}}$ $\text{7.33}^{\text{{\color[rgb]{0.68,0,0}-33.88}}}$ $\text{9.21}^{\text{{\color[rgb]{0.68,0,0}-24.77}}}$ $\text{7.76}^{\text{{\color[rgb]{0.68,0,0}-35.05}}}$
4Docc $-28.68\text{8.54}^{\text{{\color[rgb]{0.68,0,0}{-28.68}}}}$ $-27.87\text{13.35}^{\text{{\color[rgb]{0.68,0,0}{-27.87}}}}$ $-25.82\text{8.15}^{\text{{\color[rgb]{0.68,0,0}{-25.82}}}}$ $\text{13.45}^{\text{{\color[rgb]{0.68,0,0}-29.36}}}$
Uno $-0.57\text{36.66}^{\text{{\color[rgb]{0.68,0,0}{-0.57}}}}$ $-6.19\text{35.03}^{\text{{\color[rgb]{0.68,0,0}{-6.19}}}}$ $-6.15\text{27.83}^{\text{{\color[rgb]{0.68,0,0}{-6.15}}}}$ $\text{37.41}^{\text{{\color[rgb]{0.68,0,0}-5.41}}}$
Ours $\textbf{44.68}^{\textbf{{\color[rgb]{0,0.88,0}+7.46}}}$ $\textbf{43.29}^{\textbf{{\color[rgb]{0,0.88,0}+2.07}}}$ $\textbf{43.36}^{\textbf{{\color[rgb]{0,0.88,0}+9.39}}}$ $\textbf{49.37}^{\textbf{{\color[rgb]{0,0.88,0}+6.55}}}$
20% No 47.41 47.38 45.39 48.77
ALSO $\text{23.37}^{\text{{\color[rgb]{0.68,0,0}-24.04}}}$ $\text{23.02}^{\text{{\color[rgb]{0.68,0,0}-24.36}}}$ $\text{23.37}^{\text{{\color[rgb]{0.68,0,0}-22.02}}}$ $\text{23.02}^{\text{{\color[rgb]{0.68,0,0}-25.75}}}$
4Docc $-21.80\text{25.61}^{\text{{\color[rgb]{0.68,0,0}{-21.80}}}}$ $-11.55\text{35.83}^{\text{{\color[rgb]{0.68,0,0}{-11.55}}}}$ $-20.16\text{25.23}^{\text{{\color[rgb]{0.68,0,0}{-20.16}}}}$ $\text{36.24}^{\text{{\color[rgb]{0.68,0,0}-12.53}}}$
Uno $\text{45.48}^{\text{{\color[rgb]{0.68,0,0}-1.92}}}$ $\text{42.09}^{\text{{\color[rgb]{0.68,0,0}-5.29}}}$ $\text{43.77}^{\text{{\color[rgb]{0.68,0,0}-1.62}}}$ $\text{46.09}^{\text{{\color[rgb]{0.68,0,0}-2.68}}}$
Ours $\textbf{58.45}^{\textbf{{\color[rgb]{0,0.88,0}+11.04}}}$ $\textbf{50.21}^{\textbf{{\color[rgb]{0,0.88,0}+2.82}}}$ $\textbf{58.45}^{\textbf{{\color[rgb]{0,0.88,0}+13.06}}}$ $\textbf{50.21}^{\textbf{{\color[rgb]{0,0.88,0}+1.44}}}$
50% No 53.13 56.35 48.20 59.20
ALSO $\text{49.82}^{\text{{\color[rgb]{0.68,0,0}-3.32}}}$ $\text{57.07}^{\text{{\color[rgb]{0,0.88,0}+0.72}}}$ $\text{44.27}^{\text{{\color[rgb]{0.68,0,0}-3.93}}}$ $\text{60.25}^{\text{{\color[rgb]{0,0.88,0}+1.04}}}$
4Docc $\text{39.23}^{\text{{\color[rgb]{0.68,0,0}-13.91}}}$ $\text{44.48}^{\text{{\color[rgb]{0.68,0,0}-11.88}}}$ $\text{38.36}^{\text{{\color[rgb]{0.68,0,0}-9.84}}}$ $\text{44.87}^{\text{{\color[rgb]{0.68,0,0}-14.33}}}$
Uno $\text{53.61}^{\text{{\color[rgb]{0,0.88,0}+0.48}}}$ $\text{54.97}^{\text{{\color[rgb]{0.68,0,0}-1.38}}}$ $\textbf{52.73}^{\textbf{{\color[rgb]{0,0.88,0}+4.53}}}$ $\text{57.06}^{\text{{\color[rgb]{0.68,0,0}-2.14}}}$
Ours $\textbf{55.81}^{\textbf{{\color[rgb]{0,0.88,0}+2.67}}}$ $\textbf{58.85}^{\textbf{{\color[rgb]{0,0.88,0}+2.50}}}$ $\text{52.11}^{\text{{\color[rgb]{0,0.88,0}+3.91}}}$ $\textbf{61.55}^{\textbf{{\color[rgb]{0,0.88,0}+2.35}}}$
Table 2: Cross-dataset transferring results on the SemanticKITTI MOS benchmark. The best performance in each column is in bold. The metric used for model selection is in bold in the table header.
5.4 Semantic Segmentation Results on nuScenes
We also assess how well our pre-training generalizes to a new downstream task by conducting a few-shot semantic segmentation experiment on the nuScenes dataset. We fine-tune the pre-trained encoders for a sufficient number of epochs using 10% of the train split sequences. As shown in Tab. 3, TOP shows a +0.42% mIoU improvement compared to the training-from-scratch baseline, while the performance of other methods notably declines. We suspect that the performance drop of 4Docc and UnO may be due to the complex “flow” information causing network overfitting, as discussed in Sec. 1. This experiment demonstrates the potential generalization capability of TOP pre-training across different downstream tasks.
Pretrain No ALSO 4Docc UnO Ours
mIoU 50.55 $50.26^{\textbf{{\color[rgb]{0.68,0,0}-0.29}}}$ $47.49^{\textbf{{\color[rgb]{0.68,0,0}-3.06}}}$ $48.06^{\textbf{{\color[rgb]{0.68,0,0}-2.49}}}$ $\textbf{50.97}^{\textbf{{\color[rgb]{0,0.88,0}+0.42}}}$
Table 3: Few-shot results on nuScenes semantic segmentation.
5.5 Ablation Study on Training Objectives
We perform an ablation study on the two training objectives $\mathcal{L}_{\text{overlap}}$ and $\mathcal{L}_{\text{recon}}$. Results in Tab. 4 indicate that employing both objectives enhances overall performance compared to using only the temporal overlapping prediction objective. As current occupancy reconstruction is straightforward and its loss converges rapidly, we speculate it acts as a regularization term to prevent overfitting.
Objectives Best $\textbf{mIoU}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
$\textbf{mIoU}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{mIoU}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
$\mathcal{L}_{\text{overlap}}$ 24.559 34.785 22.639 35.350
$\mathcal{L}_{\text{overlap}}+\mathcal{L}_{\text{recon}}$ 25.75 35.94 25.75 35.94
Table 4: Ablation study on two training objectives.
Figure 4: Visualization of the nuScenes MOS results.
Figure 5: Visualization of the SemanticKITTI MOS results.
6 Conclusion
In this paper, we propose a self-supervised pre-training method called Temporal Overlapping Prediction (TOP) for moving object segmentation (MOS). TOP predicts the occupancy states of temporal overlapping points to help the model learn spatiotemporal representations for MOS. We evaluate the effectiveness of TOP through few-shot experiments on nuScenes and cross-dataset transfer experiments on SemanticKITTI. TOP consistently improves both the object-level metric and the point-level metric, demonstrating strong transferability across LiDAR setups. Additionally, the improved performance in semantic segmentation suggests potential generalization to more downstream tasks. Further investigation of the effectiveness of TOP in other temporal tasks such as object tracking is worthwhile.

### v2

@@
 Abstract
-Moving object segmentation (MOS) on LiDAR point clouds is crucial for autonomous systems like self-driving vehicles.
-Previous supervised approaches rely heavily on costly manual annotations, while LiDAR sequences naturally capture temporal motion cues that can be leveraged for self-supervised learning.
-In this paper, we propose Temporal Overlapping Prediction (TOP), a self-supervised pre-training method that alleviate the labeling burden for MOS.
-TOP explores the temporal overlapping points that commonly observed by current and adjacent scans, and learns spatiotemporal representations by predicting the occupancy states of temporal overlapping points.
-Moreover, we utilize current occupancy reconstruction as an auxiliary pre-training objective, which enhances the current structural awareness of the model.
-We conduct extensive experiments and observe that the conventional metric Intersection-over-Union (IoU) shows strong bias to objects with more scanned points, which might neglect small or distant objects.
-To compensate for this bias, we introduce an additional metric called $\text{mIoU}_{\text{obj}}$ to evaluate object-level performance.
-Experiments on nuScenes and SemanticKITTI show that TOP outperforms both supervised training-from-scratch baseline and other self-supervised pre-training baselines by up to 28.77% relative improvement, demonstrating strong transferability across LiDAR setups and generalization to other tasks.
-Code and pre-trained models will be publicly available upon publication.
+Moving object segmentation (MOS) on LiDAR point clouds is crucial for autonomous systems such as self-driving vehicles.
+While previous supervised approaches rely on costly manual annotations, LiDAR sequences naturally capture temporal motion cues that can be leveraged for self-supervised learning.
+In this paper, we propose Temporal Overlapping Prediction (TOP), a self-supervised pre-training method designed to alleviate this annotation burden.
+TOP learns powerful spatiotemporal representations by predicting the occupancy states of temporal overlapping points that are commonly observed in current and adjacent scans.
+To further ground these representations in the current scene’s geometry, we introduce an auxiliary pre-training objective of reconstructing the occupancy of the current scan.
+Extensive experiments on the nuScenes and SemanticKITTI datasets validate our method’s effectiveness.
+TOP consistently outperforms existing supervised and self-supervised pre-training baselines across both point-level Intersection-over-Union (IoU) and object-level Recall metrics.
+Notably, it achieves a relative improvement of up to 28.77% over a training-from-scratch baseline and demonstrates strong transferability across LiDAR setups.
+Our code is publicly available at https://github.com/ZiliangMiao/TOP.
 1 Introduction
-Moving Object Segmentation (MOS) on LiDAR data aims to segment moving points from the entire LiDAR scan and provides autonomous systems [9, 33, 34, 50, 54] with important clues in localization and dynamic obstacle avoidance, which enables better planning of collision-free trajectories in dynamic scenes.
+Figure 1: Temporal occupancy inconsistency caused by the relative motion between the sensor and the person.
+The red dotted lines track fixed points (temporal overlapping points) in space across time, illustrating how their occupancy states change.
+The point colors denote the occupancy state, while the index indicates the observation time.
+Moving Object Segmentation (MOS) on LiDAR data aims to segment moving points from the entire LiDAR scan, providing autonomous systems [9, 33, 34, 50, 55] with important clues for localization and dynamic obstacle avoidance, thereby enabling better planning of collision-free trajectories in dynamic scenes.
 However, per-point manual labeling of moving objects [9, 3, 6] is both costly and labor-intensive.
-Meanwhile, self-supervised pre-training achieves significant success in Natural Language Processing (NLP) [13, 30, 19, 40, 5], Computer Vision (CV) [51, 17, 18, 14, 41], and 3D Vision [59, 39, 55, 58], enabling the learning of transferable embeddings from unlabeled raw data.
-This motivates us to pre-train a backbone using raw LiDAR data in a self-supervised manner and then fine-tune it for MOS, which has limited expensive labeled data.
-The MOS task heavily relies on temporal information, including the positions of an object at different times and the changing relative positions of an object to its surroundings.
+Meanwhile, self-supervised pre-training has achieved significant success in natural language processing [13, 30, 19, 40, 5], computer vision [51, 17, 18, 14, 41], and 3D vision [60, 39, 56, 59], enabling the learning of transferable embeddings from unlabeled raw data.
+This motivates us to pre-train a backbone using raw LiDAR data in a self-supervised manner and then fine-tune it for MOS, where labeled data is limited and expensive.
+The MOS task heavily relies on temporal information to discern the motion of objects relative to their surroundings.
 In the literature involving temporal information of LiDAR data, an intuitive approach is to utilize LiDAR observation forecasting as a self-supervised pre-training method for MOS.
@@
 For moving objects, beyond geometry, it is essential to extract temporal features such as the object’s kinematics and interactions between objects and their surroundings.
-Hence, relying on forecasting as a pre-training step for MOS could be overly demanding.
+Hence, relying on forecasting as a pre-training step for MOS could be unnecessarily complex.
 (2) Existing methods [23, 1] inevitably learn a “flow” when using raw future observations as supervision.
-This complex “flow” information represents the relationship between past observations’ embeddings and future observations.
+This complex “flow” information represents the relationship between the embeddings of past observations and future observations.
 Current approaches either implicitly learn this flow through simple feature broadcasting [23] or explicitly predict it [1], lacking direct supervision similar to the scene flow task [62, 63, 29].
 Thus, the learned “flow” tends to be noisy.
-Relying on this noisy “flow” for deterministic future observation predictions (as opposed to the generative approach [60]) can lead to network overfitting, thereby reducing its transferability to target downstream tasks.
-Figure 1: Illustration of temporal overlapping points in a realistic scenario.
-The color of points represents the occupancy state, and the index marked on the points indicates the observed time.
-For the two temporal overlapping points in the past and one in the future, their temporal occupancy states differ from the current occupancy states.
-To overcome these two limitations and incorporate temporal information for MOS self-supervised pre-training, we propose Temporal Overlapping Prediction.
-TOP explores the temporal occupancy states of temporal overlapping points observed by both the current scan and adjacent past or future scans for pre-training.
-The rationale for this lies in the potential for moving objects to cause the current occupancy states of temporal overlapping points to differ from the states at other observed times, as illustrated in Figure 1.
-We first pre-process the spatial positions and temporal occupancy states of the overlapping points with beam divergence [31], which is a physical characteristic of LiDAR.
-This LiDAR nature helps avoid discretization errors caused by manually setting a fixed voxel resolution.
-Then, we pre-train the encoder by predicting the pre-processed temporal occupancy states, which simplifies the task compared to directly forecasting LiDAR observations.
-Based on a distinct temporal correlation between LiDAR beams at different times established by the pre-processing, our pre-training factors out either implicit or explicit learning of a noisy “flow”.
-Additionally, since understanding current states is also important for MOS, we add a reconstruction term for optimization.
-We assess the MOS performance on the nuScenes [6] and SemanticKITTI [3] datasets.
-Previous evaluations of nuScenes MOS in [34, 46] directly convert object attributes to MOS labels.
-However, these attributes do not precisely reflect the motion state of the object, and not all objects have attributes.
-Thus, we combine object attributes with object speed calculated based on bounding boxes to obtain refined nuScenes MOS labels.
-In the experiments, we observe that the conventional metric Intersection-over-Union (IoU) for MOS exhibits a significant bias towards objects that have a larger number of scanned points, potentially overlooking smaller or more distant objects.
-Thus, we propose an additional object-level evaluation metric $\text{mIoU}_{\text{obj}}$ to take these objects into account.
-We conducted few-shot fine-tuning experiments on the nuScenes dataset with different amounts of labeled data.
-Results show that TOP achieves improvements over the supervised training-from-scratch baseline [33] and outperforms other self-supervised pre-training baselines, such as state-of-the-art (SOTA) LiDAR observation forecasting methods [23, 1] and the occupancy-based 3D pre-training method [4].
-The results of our few-shot cross-dataset experiments on the SemanticKITTI dataset further highlight the superior transferability of TOP pre-training compared to various baselines.
-Additionally, we fine-tune the pre-trained backbone on the nuScenes semantic segmentation task in a few-shot setting, demonstrating the potential generalization capability of TOP across more tasks.
+Relying on this noisy “flow” for deterministic future observation predictions (as opposed to the generative approach [61]) can lead to network overfitting, thereby reducing its transferability to the downstream task.
+To overcome these two limitations and incorporate temporal information for MOS self-supervised pre-training, we propose TOP, namely Temporal Overlapping Prediction.
+TOP explores the occupancy states of temporal overlapping points observed by both the current scan and adjacent (past or future) scans for pre-training.
+The rationale is that the motion of the sensor and objects creates temporal occupancy inconsistency, where a temporal overlapping point may exhibit different occupancy states when observed at different times, as shown in Fig 1.
+We first pre-process the spatial positions and occupancy states of the temporal overlapping points by considering beam divergence [31], which is a physical characteristic of LiDAR.
+Leveraging this physical characteristic of LiDAR helps avoid discretization errors caused by manually setting a fixed voxel resolution.
+Then, we pre-train the encoder by predicting the pre-processed occupancy states, which simplifies the task compared to directly forecasting LiDAR observations.
+Based on a distinct temporal correlation between LiDAR beams at different times established by the pre-processing, our pre-training avoids the need for either implicitly or explicitly learning a noisy “flow”.
+Additionally, since understanding the current scene geometry is also important for MOS, we include a reconstruction term in our pre-training objective.
+We evaluate TOP via few-shot fine-tuning on nuScenes [6] with varying amounts of labeled data.
+The results demonstrate that TOP outperforms both the supervised training-from-scratch baseline [33] and other self-supervised pre-training baselines, including state-of-the-art (SOTA) LiDAR forecasting methods [23, 1] and the occupancy-based 3D method [4].
+Furthermore, cross-dataset experiments on SemanticKITTI [3] highlight TOP’s superior transferability, while its few-shot performance on the nuScenes semantic segmentation task demonstrates its potential to generalize to other downstream tasks.
 2 Related Works
@@
 Existing supervised MOS methods can be grouped into projection-based and voxel-based approaches.
-The early MOS methods [9, 24, 45, 11] project 3D LiDAR data to range images and embed the range residual images by 2D backbone.
+The early MOS methods [9, 24, 45, 11] project 3D LiDAR data onto range images and embed the range residual images using a 2D backbone.
 These methods suffer from imprecise boundaries when re-projecting 2D outputs to 3D point clouds [34].
-Recent methods [46, 33, 34, 50] utilize the sparse convolution backbone to directly embed voxelized point clouds.
-4DMOS [33] first employs sparse 4D UNet [12] and makes it the SOTA backbone for MOS.
-For unsupervised approaches, several scene flow methods [48, 2, 16, 44] decouple the per-point flow to an ego-motion flow for static points and a non-rigid flow for moving points, which provides a byproduct of MOS [33].
-However, these approaches using only two adjacent scans, encounter challenges in identifying slowly moving objects [33].
-MOTS [25] proposes a label-free method that only works in stationary scenes.
-4dNDF [64] encodes the 4D scene into an implicit neural map, which can be used to distinguish the static map and dynamic objects.
-However, its application in real-time is limited.
+Recent methods [46, 33, 34, 50] utilize a sparse convolution backbone to directly embed voxelized point clouds.
+4DMOS [33] first employed sparse 4D UNet [12] and established it as the SOTA backbone for MOS.
+For unsupervised approaches, several scene flow methods [48, 2, 16, 44] decouple the per-point flow into an ego-motion flow for static points and a non-rigid flow for moving points, which yields MOS as a byproduct.
+However, these two-frame approaches encounter challenges in identifying slowly moving objects [33].
+MOTS [25] proposes a label-free method that is limited to stationary scenes.
+4dNDF [65] encodes the 4D scene into an implicit neural map, which can be used to distinguish between the static map and dynamic objects.
+However, its computational cost limits real-time applicability.
 In addition, M-detector [55] proposes a network-free approach based on the occlusion principle.
-It requires manual tuning of numerous hyper-parameters when using different LiDARs or deploying in varying scenes.
-Moreover, to reduce the cost of manual labeling, several methods [10, 28] propose a pipeline that first generates coarse labels automatically and then refines them manually.
-In contrast to previous works on MOS, we employ self-supervised pre-training to enhance the few-shot MOS performance.
+It requires manual tuning of numerous hyperparameters when using different LiDARs or deploying in different scenes.
+To reduce manual labeling costs, several methods [10, 28] propose automatic labeling pipelines, though some of them still incorporate a manual refinement step.
+In contrast to previous MOS works, we employ self-supervised pre-training to enhance the few-shot MOS performance.
 3D Self-supervised Pre-training.
-Effective self-supervised pre-training methods in NLP [13, 40, 5] and CV [18] have been extended to 3D data.
+Effective self-supervised pre-training methods in natural language processing [13, 40, 5] and computer vision [18] have been extended to 3D data.
 Some methods [60, 15, 7] leverage BERT-like [13] or GPT-like [40, 5] pre-training on object CAD point clouds.
-Contrastive learning methods [55, 42, 38, 22, 27, 8] employ various transformations or augmentations to construct point/instance/scene-level contrastive loss, learning geometric or semantic information without manual labels.
-Additionally, Masked Auto-Encoder (MAE)-based methods [39, 63, 20, 57, 47, 36, 26, 56] extend MAE [18] to 3D point clouds.
-These methods train an encoder by predicting a certain ratio of masked point cloud patches, which transfers well on object detection and semantic segmentation tasks.
-Subsequent works [21, 65, 58] apply NeRF-like [35] volumetric neural rendering to leverage multi-modal inputs and bridge the gap between 2D and 3D data.
-In summary, existing 3D self-supervised pre-training works lack temporal information and thus are unsuitable for MOS.
+Contrastive learning methods [56, 42, 22, 27, 8] employ various transformations or augmentations for contrastive losses at point, instance, and scene levels, learning geometric or semantic information without manual labels.
+Inspired by MAE [18], numerous methods [39, 64, 20, 58, 47, 36, 26, 57] pre-train an encoder by reconstructing masked point cloud patches, which yields effective 3D representations for object detection and semantic segmentation.
+Subsequent works [21, 66, 59] apply NeRF-like [35] volumetric neural rendering to integrate 2D and 3D multi-modal data.
+To the best of our knowledge, while a few methods [38, 52] incorporate temporal information, TOP is the first self-supervised pre-training method specifically designed for dynamic tasks like MOS.
 LiDAR Observation Forecasting.
 Existing works that involve temporal information of LiDAR data mainly focus on LiDAR observation forecasting, including point cloud and occupancy forecasting.
-Previous works on point cloud forecasting [53, 52, 32] directly predict the raw future LiDAR scans based on past ones.
-A recent study, 4Docc [23], claims the potential of forecasting as a scalable self-supervised task.
+Previous works on point cloud forecasting [54, 53, 32] directly predict the future LiDAR scans based on past ones.
+A recent study, 4DOcc [23], highlights the potential of forecasting as a scalable self-supervised task.
 Instead of directly predicting raw point clouds, 4DOcc renders the future point clouds on a predicted occupancy map using known sensor intrinsics and extrinsics, enabling the network to focus on learning the world rather than sensor characteristics.
-Another generative approach, Copilot4D [60] introduces the concept of the world model.
+Another generative approach, Copilot4D [61], introduces the concept of the world model.
 It tokenizes LiDAR scans and predicts future scans through discrete diffusion.
-However, the learned world knowledge is difficult to apply to downstream tasks due to the nature of heavy-decoder.
-ALSO [4] is a single-scan method that reconstructs the occupancy of a local surface and inspires UnO [1].
-UnO is the most recent occupancy forecasting method that predicts the binary occupancy class of points sampled along future LiDAR rays.
-In this paper, we pre-train the backbone using the proposed TOP method instead of future observation forecasting.
+However, the learned world knowledge is difficult to apply to downstream tasks due to its heavy-decoder architecture.
+ALSO [4] is a single-scan method that reconstructs the occupancy of a local surface.
+This work inspired UnO [1], which is the most recent occupancy forecasting method that predicts the binary occupancy class of points sampled along future LiDAR rays.
+In contrast, our work diverges from these forecasting-based approaches by pre-training the backbone with our proposed TOP method.
 3 Methodology
-Figure 2: The overall pipeline.
-Pre-processing: (a) Coplanarity condition of two beams.
+Figure 2: LiDAR beam divergence: The diverging beam is illustrated as the blue cone.
+The reported point (black) lies on the beam’s centerline, while the actual beam hit-point (red) can be anywhere within the beam’s footprint.
+LiDAR occupancy measurement: We model the space from the sensor to the reported LiDAR point as free with maximum confidence.
+Beyond this point, confidence decays exponentially.
+The region immediately following the reported point is considered occupied if its confidence is above a predefined threshold.
+The subsequent region where the confidence falls below the threshold is considered unknown.
+Figure 3: Overall pipeline.
+Pre-processing: (a) shows the coplanarity condition of two beams.
 (b) If the spatial angle between two beams $\alpha_{i,j}$ exceeds the beam divergence angle $\theta_{\text{dvg}}$, we sample the beams’ intersection point as the temporal overlapping point.
 The red and blue areas indicate beam divergence.
-(c) When $\alpha_{i,j}$ is less than $\theta_{\text{dvg}}$, the intersection segment is calculated to sample temporal overlapping points within it.
-Pre-training: (d) We use a sparse 4D UNet to embed the input scans.
-For the temporal overlapping prediction objective $\mathcal{L}_{\text{overlap}}$, the occupancy state of a temporal overlapping point is predicted based on its positional encoding and the embedding of the current beam hit-point.
-Similarly, the occupancy state of the currently sampled point is also predicted for the current reconstruction objective $\mathcal{L}_{\text{recon}}$.
-In this section, we introduce the self-supervised pre-training method TOP.
-Our method aims to pre-train an encoder for the downstream MOS task.
-The pre-training goal is to predict the occupancy states of temporal overlapping points and to reconstruct the current occupancy.
-We first introduce notations and problem formulation in Sec.
+(c) When $\alpha_{i,j}$ is less than $\theta_{\text{dvg}}$, we calculate the intersection segment and sample temporal overlapping points from it.
+Pre-training: (d) The input sequence is encoded by a sparse 4D UNet.
+For both pre-training objectives, a shallow MLP decoder predicts occupancy states based on the point’s positional encoding and the feature of its corresponding LiDAR point.
+In this section, we introduce TOP, a self-supervised method to pre-train an encoder for the downstream MOS task.
+The pre-training goal involves two objectives: predicting the occupancy of temporal overlapping points and reconstructing the current scene’s occupancy.
+We first introduce the notation and problem formulation in Sec.
 3.1.
@@
 To begin with, we define a LiDAR scan as $\mathbf{P}\in\mathbb{R}^{N\times 3}$, where $N$ is the number of points.
-We use subscript $i$ to indicate the $i^{th}$ point in $\mathbf{P}$, where each point $\mathbf{p}_{i}\in\mathbb{R}^{3}$ ($i=1,2,...,N$) contains $xyz$ location.
-$\mathbf{p}_{i}$ can be further described as a laser beam emitted from the sensor origin $\mathbf{a}\in\mathbb{R}^{3}$ with its normalized beam direction $\mathbf{d}_{i}\in\mathbb{R}^{3}$ and range $r_{i}$, leading to $\mathbf{p}_{i}=\mathbf{a}+r_{i}\mathbf{d}_{i}$.
-Although the laser beam is commonly modeled as an ideal ray, the real LiDAR beam exhibits a diverging radius as it travels, and this is referred to as beam divergence [31].
+We use subscript $i$ to indicate the $i^{th}$ point in $\mathbf{P}$, where each point $\mathbf{p}_{i}\in\mathbb{R}^{3}$ ($i=1,2,\dots,N$) contains xyzxyz coordinates.
+$\mathbf{p}_{i}$ can be further described by a laser beam emitted from the sensor origin $\mathbf{a}\in\mathbb{R}^{3}$ with its normalized beam direction $\mathbf{d}_{i}\in\mathbb{R}^{3}$ and range $r_{i}$, leading to $\mathbf{p}_{i}=\mathbf{a}+r_{i}\mathbf{d}_{i}$.
+Although the laser beam is commonly modeled as an ideal ray, an actual LiDAR beam exhibits a diverging radius as it travels, a phenomenon known as beam divergence [31].
 The beam divergence angle, denoted as $\theta_{\text{dvg}}$, is a constant for each type of LiDAR.
-When any part of the diverging beam hits an object, the LiDAR calculates the distance by time of flight and considers the beam center point as the hit-point, as shown in Fig.
-3.
+When a diverging LiDAR beam hits an obstacle and reflects, the sensor reports a point on the beam’s centerline corresponding to the measured time-of-flight range.
+In contrast, the actual beam hit-point (reflection point) can be located anywhere within the beam’s footprint, as shown in Fig.
+2.
 To incorporate the temporal LiDAR sequence, we indicate scans at different timestamps with superscripts.
-For example, the scan at the current timestamp $t_{0}$ is $\mathbf{P}^{t_{0}}\in\mathbb{R}^{N\times 3}$.
-For adjacent past/future timestamps $\mathcal{T}=\{t_{-n}\ldots t_{-1},t_{1}\ldots t_{n}\}$, we have $\mathcal{P}_{\text{adj}}=\{\mathbf{P}^{t}\in\mathbb{R}^{N_{t}\times 3}\}_{t\in\mathcal{T}}$.
-The $j^{th}$ beam at time $t$, denoted as $\mathbf{d}_{j}^{t}$, can also be described by $\mathbf{p}_{j}^{t}=\mathbf{a}^{t}+r_{j}^{t}\mathbf{d}_{j}^{t}$.
-Point Clouds Encoding.
-All scans in $\mathcal{P}_{\text{adj}}$ are transformed to the local coordinate system of $\mathbf{P}^{t_{0}}$ with the poses provided in the datasets.
-We take $n$ current and past scans as network inputs $\mathcal{P}_{\text{in}}=\{\mathbf{P}^{t_{-n+1}}\ldots\mathbf{P}^{t_{-1}},\mathbf{P}^{t_{0}}\}$.
-Each point in $\mathcal{P}_{\text{in}}$ is concatenated with its timestamp as a 4D point before being fed into the network.
+For example, the scan at the current timestamp $t_{0}$ is $\mathbf{P}^{t_{0}}\in\mathbb{R}^{N_{t_{0}}\times 3}$.
+For adjacent past/future timestamps $\mathcal{T}=\{t_{-n},\dots,t_{-1},t_{1},\dots,t_{n}\}$, we have $\mathcal{P}_{\text{adj}}=\{\mathbf{P}^{t}\in\mathbb{R}^{N_{t}\times 3}\}_{t\in\mathcal{T}}$.
+The $j^{th}$ beam at time $t$, $\mathbf{d}_{j}^{t}$, results in the point $\mathbf{p}_{j}^{t}$, which is described by $\mathbf{p}_{j}^{t}=\mathbf{a}^{t}+r_{j}^{t}\mathbf{d}_{j}^{t}$.
+Point Cloud Encoding.
+All scans in $\mathcal{P}_{\text{adj}}$ are transformed to the local coordinate system of $\mathbf{P}^{t_{0}}$ using the poses provided in the datasets.
+We take $n$ current and past scans as network inputs $\mathcal{P}_{\text{in}}=\{\mathbf{P}^{t_{-n+1}},\dots,\mathbf{P}^{t_{-1}},\mathbf{P}^{t_{0}}\}$.
+Each point in $\mathcal{P}_{\text{in}}$ is concatenated with its timestamp before being fed into the network.
 We apply a sparse 4D UNet encoder $g(\cdot)$ used in [12, 33, 34, 50] to embed the spatiotemporal information of the 4D inputs:
 $\mathbf{F}=g(\mathcal{P}_{\text{in}}),$ (1)
-where $\mathbf{F}\in\mathbb{R}^{N\times d}$ is the fully sparse features of the current scan $\mathbf{P}^{t_{0}}$ with a feature dimension of $d$.
+where $\mathbf{F}\in\mathbb{R}^{N_{t_{0}}\times d}$ is a sparse feature for points in the current scan $\mathbf{P}^{t_{0}}$ with a feature dimension of $d$.
 Pre-training Loss.
@@
 As described in Fig.
-1, temporal overlapping points are located on current LiDAR beams in $\mathbf{P}^{t_{0}}$ and are also observed by adjacent past/future beams in $\mathcal{P}_{\text{adj}}$.
-Firstly, we denote the temporal overlapping points as $\mathbf{O}\in\mathbb{R}^{M\times 4}$, where $M$ is the number of points and each point contains $xyz$ position and the temporally observed timestamp $t$.
-Then, we use the LiDAR occupancy measurement model (introduced in Sec.
-3.2) to describe the occupancy state labels of $\mathbf{O}$, denoted as $\mathbf{S}\in\mathbb{R}^{M\times 3}$.
-Each label is a three-dimensional one-hot vector indicating the occupancy states “free”, “occupied”, and “unknown”.
-The overall pre-training goal consists of temporal occupancy prediction and current occupancy reconstruction, which is described as:
+1, temporal overlapping points are locations along the LiDAR beams from the current scan $\mathbf{P}^{t_{0}}$ that are also intersected by beams from adjacent past/future scans $\mathcal{P}_{\text{adj}}$.
+Firstly, we denote the temporal overlapping points as $\mathbf{O}\in\mathbb{R}^{M\times 4}$, where $M$ is the number of points and each point has xyzxyz coordinates and the timestamp $t$ of the adjacent scan that observed it.
+Then, we use the LiDAR occupancy measurement (introduced in Sec.
+3.2) to determine the occupancy state labels for $\mathbf{O}$, denoted as $\mathbf{S}\in\mathbb{R}^{M\times 3}$.
+The overall pre-training loss $\mathcal{L}$, further described in Sec.
+3.3, consists of temporal overlapping prediction $\mathcal{L}_{\text{overlap}}$ and current occupancy reconstruction $\mathcal{L}_{\text{recon}}$:
 $\mathcal{L}=\mathcal{L}_{\text{overlap}}(\mathbf{O},\mathbf{S},\mathbf{F})+\mathcal{L}_{\text{recon}}(\mathbf{P}^{t_{0}},\mathbf{F}).$ (2)
-Figure 3: LiDAR beam divergence model: The bottom-left blue cone illustrates a real diverging beam, where the beam radius expands as the traversed distance increases.
-When any point within the beam spot hits an obstacle (indicated by the real hit-point colored in red), the center point of the beam spot (colored in black) is considered as the measured point.
-LiDAR occupancy measurement model: The occupancy state is confidently classified as free before reaching the beam hit-point.
-Then, the confidence level exponentially decays from the beam hit-point onward.
-A small area of high confidence after the beam hit-point is categorized as the occupied class, while the remaining areas are deemed unknown.
 3.2 Temporal Overlapping Pre-processing
-As computing temporal overlapping during pre-training can be inefficient, we pre-process the LiDAR scans $\mathbf{P}^{t_{0}}$ and $\mathcal{P}_{\text{adj}}$ to produce temporal overlapping points $\mathbf{O}$ and their occupancy states $\mathbf{S}$.
-We first identify coplanar beam pairs from $\mathbf{P}^{t_{0}}$ and $\mathcal{P}_{\text{adj}}$, as coplanarity is a necessary condition for beams to have overlapping observations, as illustrated in Fig.
-2(a).
+As computing temporal overlapping points during pre-training can be inefficient, we pre-process the LiDAR scans $\mathbf{P}^{t_{0}}$ and $\mathcal{P}_{\text{adj}}$ to produce temporal overlapping points $\mathbf{O}$ and their occupancy states $\mathbf{S}$.
+We first identify coplanar beam pairs from $\mathbf{P}^{t_{0}}$ and $\mathcal{P}_{\text{adj}}$, as coplanarity is a necessary condition for beams to overlap, as illustrated in Fig.
+3(a).
 Then, we compute the intersection of coplanar beam pairs and sample temporal overlapping points.
-Depending on the spatial angle between two beams, we split the computation and sampling into two scenarios, shown in Fig.
-2(b) and Fig.
-2(c), respectively.
-We take an example beam pair for a clearer explanation: the $i^{th}$ beam at time $t_{0}$ and the $j^{th}$ beam at time $t$ ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$), emitted from ($\mathbf{a}^{t_{0}}$, $\mathbf{a}^{t}$) and hitting ($\mathbf{p}_{i}^{t_{0}}$, $\mathbf{p}_{j}^{t}$).
+We handle the computation and sampling in two distinct scenarios based on the spatial angle between two beams, as shown in Fig.
+3(b) and Fig.
+3(c), respectively.
+For clarity, we consider a beam pair: the $i^{th}$ beam at time $t_{0}$ and the $j^{th}$ beam at time $t$ ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$), emitted from ($\mathbf{a}^{t_{0}}$, $\mathbf{a}^{t}$) and hitting ($\mathbf{p}_{i}^{t_{0}}$, $\mathbf{p}_{j}^{t}$).
 When a beam pair ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$) generates the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$, we record the $(i,j,k)$ correspondence for subsequent usage in Sec.
@@
 We define a reference plane formed by ($\mathbf{a}^{t_{0}}$, $\mathbf{a}^{t}$, $\mathbf{p}_{i}^{t_{0}}$), illustrated as the yellow triangle in Fig.
-2(a).
+3(a).
 The unit normal vector of this plane is:
-$\mathbf{n}_{i}^{t}=\mathbf{d}_{i}^{t_{0}}\times\frac{\mathbf{a}^{t}}{||\mathbf{a}^{t}||}.$ (3)
-The space angle between beam $\mathbf{d}_{j}^{t}$ and the plane is:
+[…]
+The spatial angle between beam $\mathbf{d}_{j}^{t}$ and the plane is:
 $\mathbf{\theta}_{i,j}=\arccos{(\mathbf{n}_{i}^{t}\cdot\mathbf{d}_{j}^{t})}-\frac{\pi}{2}.$ (4)
@@
 This scenario describes most overlapping cases where the spatial angle between two beams exceeds the beam divergence angle, as shown in Fig.
-2(b).
-The beam intersection is small, thus being considered as a point $\mathbf{q}_{i,j}$, which is the intersection of the two beams’ centerlines.
-The $\mathbf{q}_{i,j}$ is solved from the parameter equations of the two beams:
+3(b).
+The beam intersection is small and is thus considered a point $\mathbf{q}_{i,j}$, which is the intersection of the two beams’ centerlines.
+$\mathbf{q}_{i,j}$ is calculated from the parameter equations of the two beams:
 $\mathbf{q}_{i,j}=\frac{\mathbf{a}^{t}\times\mathbf{d}_{j}^{t}\cdot\mathbf{m}_{i,j}}{\mathbf{m}_{i,j}\cdot\mathbf{m}_{i,j}}\cdot\mathbf{d}_{i}^{t_{0}},$ (5)
-where $\mathbf{m}_{i,j}=\mathbf{d}_{i}^{t_{0}}\times\mathbf{d}_{j}^{t}$ is the unit common perpendicular line of the two beams.
+where $\mathbf{m}_{i,j}=\mathbf{d}_{i}^{t_{0}}\times\mathbf{d}_{j}^{t}$ is a vector along the common perpendicular to the two beams.
 In this scenario, the intersection point $\mathbf{q}_{i,j}$ is selected as the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$.
 Scenario-2 ($\alpha_{i,j}\leq\theta_{\text{dvg}}$).
-When the spatial angle between two beams is less than the beam divergence, the intersection of the two beams can no longer be considered as a point but an area, as shown in Fig.
-2(c).
+When the spatial angle between two beams is less than or equal to the beam divergence, the intersection of the two beams can no longer be considered a point but rather a region, as shown in Fig.
+3(c).
 In this scenario, we compute the intersection segments (purple dashed lines) on the centerlines for subsequent sampling.
 Firstly, we obtain the endpoint $\mathbf{q}_{i,j}$ from the beam intersection similar to Scenario-1.
-Then, we approximate the starting point $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ as the projection from a point $\mathbf{b}$ to the two beams.
-Here, $\mathbf{b}$ is the crossed point of the boundaries of the two beams.
+Then, we approximate the starting point $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ as the projection from a point $\mathbf{b}$ onto the two beams.
+Here, $\mathbf{b}$ is the crossing point of beam boundaries.
 Given that the spatial angle $\alpha_{i,j}$ is small, we make two approximations to compute the starting points: (1) We approximate that the two intersection segments share the same length.
-(2) We approximate that the distance between the two starting points $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ is equal to the sum of the beam radius at the two points.
-The detailed formulas for computing $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ are in the appendix.
-Since the overlap between two beams is an area, we sample 5 temporal overlapping points inside the intersection segment, as shown in Fig.
-2(c) right: $\mathbf{o}_{k_{1}}$, the beam hit point $\mathbf{p}_{i}^{t_{0}}$; $\mathbf{o}_{k_{2}}$, the projection of $\mathbf{p}_{j}^{t}$ on $\mathbf{d}_{i}^{t_{0}}$; $\mathbf{o}_{k_{3}}$, the center of ($\mathbf{o}_{k_{1}}$, $\mathbf{o}_{k_{2}}$); $\mathbf{o}_{k_{4}}$, the center of ($\mathbf{o}_{k_{1}}$, $\mathbf{q}_{i,j}$); $\mathbf{o}_{k_{5}}$, the center of ($\mathbf{o}_{k_{2}}$, $\mathbf{q}_{i,j}$).
-Their occupancy states at timestamp $t$ are given by the states of their projection points on $\mathbf{d}_{j}^{t}$.
-LiDAR Occupancy Measurement Model.
-The LiDAR occupancy measurement model [1, 43] partitions the space between the sensor origin to infinity along the beam direction into three areas with different occupancy states, as shown in Fig.
-3.
-For the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$ generated from the overlap beams pair ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$), the confidence level $w_{\text{conf}}$ of the measured occupancy state of $\mathbf{o}_{k}$ is 100% when $\mathbf{o}_{k}$ lies inside the physically traversed areas of the beam (blue area in Fig.
-3).
-It then exponentially decays after the hit-point $\mathbf{p}$, as described below:
-$w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t})=\begin{cases}1,\ \text{if}\ 0\leq||\mathbf{o}_{k}||\leq||\mathbf{p}_{j}^{t}||,\\ e^{{-(||\mathbf{o}_{k}||-||\mathbf{p}_{j}^{t}||)}},\ \text{if}\ ||\mathbf{o}_{k}||>||\mathbf{p}_{j}^{t}||.\\ \end{cases}$ (6)
-The occupancy state of $\mathbf{o}_{k}$ is denoted as $\mathbf{s}_{k}$, which is described as:
-$\mathbf{s}_{k}=\begin{cases}[1,0,0],\ \text{if}\ 0\leq||\mathbf{o}_{k}||<||\mathbf{p}_{j}^{t}||,\\ [0,1,0],\ \text{if}\ ||\mathbf{o}_{k}||=||\mathbf{p}_{j}^{t}||\ \text{or}\ \lambda_{\text{occ}}\leq w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t}),\\ [0,0,1],\ \text{if}\ w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t})<\lambda_{\text{occ}},\\ \end{cases}$ (7)
-where $\mathbf{s}_{k}$ is a one-hot vector indicating “free”, “occupied” and “unknown” occupancy states, and $\lambda_{\text{occ}}$ is the confidence level threshold of the occupied state.
+(2) We approximate that the distance between the two starting points $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ is equal to the sum of the beam radii at the two points.
+The detailed formulas for computing $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ are provided in the appendix.
+Since the overlap between two beams is a region, we sample 5 temporal overlapping points within the intersection segment, as shown in Fig.
+3(c)-right: $\mathbf{o}_{k_{1}}$, the beam hit point $\mathbf{p}_{i}^{t_{0}}$; $\mathbf{o}_{k_{2}}$, the projection of $\mathbf{p}_{j}^{t}$ on $\mathbf{d}_{i}^{t_{0}}$; $\mathbf{o}_{k_{3}}$, the midpoint of ($\mathbf{o}_{k_{1}}$, $\mathbf{o}_{k_{2}}$); $\mathbf{o}_{k_{4}}$, the midpoint of ($\mathbf{o}_{k_{1}}$, $\mathbf{q}_{i,j}$); $\mathbf{o}_{k_{5}}$, the midpoint of ($\mathbf{o}_{k_{2}}$, $\mathbf{q}_{i,j}$).
+The occupancy state for each of these sampled points at timestamp $t$ is determined by projecting it onto the adjacent beam $\mathbf{d}_{j}^{t}$ and evaluating the corresponding occupancy.
+LiDAR Occupancy Measurement.
+As shown in Fig.
+2, the LiDAR occupancy measurement [1, 43] partitions the space along each beam into three states.
+First, the space from the sensor origin to the reported point is considered free with maximum confidence.
+Beyond the reported point, an exponential confidence decay defines two subsequent states: the space is occupied while the confidence exceeds a predefined threshold, and is considered unknown where the confidence drops below the threshold.
+Given the $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$ generated from the overlapping beam pair ($\mathbf{d}_{i}^{t_{0}}$, $\mathbf{d}_{j}^{t}$), the confidence level $w_{\text{conf}}$ of $\mathbf{o}_{k}$ is formulated as:
+[…]
+The occupancy state of $\mathbf{o}_{k}$ at time $t$ is denoted as $\mathbf{s}_{k}$:
+[…]
+where $\mathbf{s}_{k}$ is a one-hot vector indicating one of three occupancy states: free, occupied, or unknown. $\lambda_{\text{occ}}$ is the predefined confidence threshold for the occupied state.
 3.3 Training Objectives
-For self-supervised pre-training, the primary objective is temporal overlapping prediction, which involves predicting $\mathbf{s}_{k}$ from the $i^{th}$ current point’s feature $\mathbf{f}_{i}$ and the positional information of $\mathbf{o}_{k}$ using a shallow MLP decoder $h(\cdot)$:
+Our primary pre-training objective, temporal overlapping prediction, is to predict the occupancy state $\mathbf{s}_{k}$ of $k^{th}$ temporal overlapping point $\mathbf{o}_{k}$.
+We predict the occupancy based on the positional encoding of $\mathbf{o}_{k}$ and the $i^{th}$ current LiDAR point’s feature $\mathbf{f}_{i}$ by a shallow MLP decoder $h(\cdot)$:
 $\mathbf{u}_{k}=h(f(\mathbf{o}_{k})+\mathbf{f}_{i})\in\mathbb{R}^{3},$ (8)
-where $\mathbf{u}_{k}$ is the predicted probabilities of three occupancy states, and $f(\mathbf{o}_{k})\in\mathbb{R}^{d}$ denotes the sine-cosine 4D positional encoding vector as described in [49].
-Temporal overlapping prediction is described as a Cross-Entropy loss:
+where $\mathbf{u}_{k}$ is the predicted probability over the three occupancy states, and $f(\mathbf{o}_{k})\in\mathbb{R}^{d}$ denotes the sine-cosine 4D positional encoding vector as described in [49].
+Temporal overlapping prediction is formulated as a Cross-Entropy loss:
 $\mathcal{L}_{\text{overlap}}=-\frac{1}{M}\sum\limits_{k=1}^{M}w_{\text{conf}}(\mathbf{o}_{k},\mathbf{p}_{j}^{t})\cdot\mathbf{w}_{s}\cdot\mathbf{s}_{k}\cdot\log{(\mathbf{u}_{k})},$ (9)
-where $\mathbf{w}_{s}$ is a constant weight vector for each occupancy state, and the $(i,j,k)$ correspondence is introduced in Sec. 3.2.
-Additionally, we reconstruct the current occupancy using the input scans, which serves as a regularization term in the total loss.
-We randomly select $M^{\prime}$ occupied and free points on each current beam.
-For the $k^{th}$ point $\mathbf{c}_{k}$ sampled on the $i^{th}$ current beam, it includes the spatial position and timestamp $t_{0}$.
+where $\mathbf{w}_{s}$ is a constant weight vector for each occupancy state, and the $(i,j,k)$ correspondence was introduced in Sec. 3.2.
+Additionally, we reconstruct the current occupancy using the input scans, which serves as a regularization term in the overall loss.
+We randomly sample $M^{\prime}$ occupied and free points along each current beam.
+For the $k^{th}$ point $\mathbf{c}_{k}$ sampled on the $i^{th}$ current beam, it consists of a spatial position and timestamp $t_{0}$.
 The ground truth occupancy label is $\mathbf{s}_{k}$, and the predicted probability is $\mathbf{v}_{k}$.
-The reconstruction objective is also specified as a Cross-Entropy loss:
+The reconstruction objective is also formulated as a Cross-Entropy loss:
 $\mathbf{v}_{k}$ $=h(f(\mathbf{c}_{k})+\mathbf{f}_{i})\in\mathbb{R}^{3},$ (10)
-$\mathcal{L}_{\text{recon}}$ $=-\frac{1}{NM^{\prime}}\sum\limits_{k=1}^{NM^{\prime}}\mathbf{w}_{s}\cdot\mathbf{s}_{k}\cdot\log{(\mathbf{v}_{k})}.$ (11)
+$\mathcal{L}_{\text{recon}}$ $=-\frac{1}{N_{t_{0}}M^{\prime}}\sum\limits_{i=1}^{N_{t_{0}}}\sum\limits_{k=1}^{M^{\prime}}\mathbf{w}_{s}\cdot\mathbf{s}_{k}\cdot\log{(\mathbf{v}_{k})}.$ (11)
 4 Evaluation Metrics
-In this section, we propose a mean object-level metric denoted as $\text{mIoU}_{\text{obj}}$ to tackle the biased weighting issue in the conventional IoU metric.
-Previous methods [9, 33, 34, 50] have relied solely on IoU as the evaluation metric.
-However, IoU implicitly gives larger weight to objects with more scanned points due to its point-wise nature.
-Objects of varying geometric sizes and distances to the LiDAR may result in different numbers of scanned points.
-Nevertheless, these objects are equally important in practical MOS applications and should not be weighted based on the number of scanned points.
-To further illustrate the impact of this bias on MOS performance evaluation, we conduct a statistical analysis: we sort all moving objects in the nuScenes train-val split based on the number of scanned points, and then calculate the cumulative distribution of moving objects and scanned points.
-The results indicate that the top 75.22% of objects with the fewest scanned points only contribute to 15.49% of the total points, while 90.06% of objects contribute to 37.02% of the points, revealing that a few objects with more points heavily influence the conventional IoU metric.
-The cumulative distribution is provided in the appendix.
-The proposed metric $\text{mIoU}_{\text{obj}}$ calculates the IoU for each moving object and computes the average, ensuring equal weighting for each moving object:
-$\text{mIoU}_{\text{obj}}=\frac{1}{m}\sum\limits_{i=1}^{m}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}},$ (12)
-where $m$ denotes the number of moving objects, $\text{TP}_{i}$ and $\text{FP}_{i}$ indicate the true positive and false positive classifications of the points belonging to $i^{th}$ moving object.
-False negatives are not involved in the $\text{mIoU}_{\text{obj}}$ as all points are classified as the moving (positive) class.
-Additionally, in the nuScenes train-val split, ego vehicles’ moving points account for 76.72% of all moving points, which can heavily influence the IoU.
-However, in autonomous driving scenarios, we already know the movement of the ego vehicle and do not need to predict it.
-Therefore, we utilize a refined IoU metric called $\text{IoU}_{\text{w/o}}$, calculated as described in [9, 33, 34] but excluding the scanned ego vehicle points.
-For a more precise and robust evaluation of the MOS performance, we use both metrics: $\text{mIoU}_{\text{obj}}$ assesses performance on all moving objects, while $\text{IoU}_{\text{w/o}}$ reflects overall performance, incorporating false negatives.
+We introduce a robust evaluation for MOS, designed to overcome two critical biases in the conventional IoU metric [9, 33, 34, 50]: the over-weighting of objects with more points, and the inflation caused by the ego vehicle’s points.
+A primary issue with the conventional IoU is its inherent bias towards objects with more points, leading to the neglect of smaller or distant objects.
+This poses significant safety risks, as all dynamic objects are equally critical for safe operation, regardless of their point count.
+For instance, a high-speed vehicle might fail to detect a distant car, or a UAV could miss a fast-approaching object.
+To quantify this imbalance, our statistical analysis of the nuScenes dataset reveals that the smallest 75.22% of objects (by point count) account for only 15.49% of the total points (see appendix).
+This confirms that a few large objects heavily dominate the final IoU score.
+The second bias is the inflation of IoU.
+In practice, the ego vehicle’s state is known via onboard sensors, and is not part of the perception challenge.
+Including these easy-to-predict points (which account for 76.72% of all moving points in the nuScenes dataset) severely inflates the metric, masking the model’s actual performance on external dynamic objects.
+To address these limitations, we use two complementary metrics.
+First, to counteract the object-size bias, we use an object-level metric $\text{Recall}_{\text{obj}}$, which ensures an equal weight of each moving object by averaging their individual recalls.
+It is computed as:
+$\text{Recall}_{\text{obj}}=\frac{1}{m}\sum\limits_{i=1}^{m}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FN}_{i}},$ (12)
+where $m$ is the number of moving objects, while $\text{TP}_{i}$ and $\text{FN}_{i}$ are the number of true positive and false negative points, respectively, within the ground truth bounding box of the $i^{th}$ object.
+Second, to resolve the ego-vehicle bias, we simply exclude all ego points from the IoU calculation and rename the metric $\text{IoU}_{\text{w/o}}$.
+This provides a more focused assessment of the model’s performance on external dynamic objects.
+Notably, for the SemanticKITTI benchmark, $\text{IoU}_{\text{w/o}}$ is equivalent to the conventional IoU, as the benchmark already omits ego points by classifying them as unlabeled.
+To demonstrate that the conventional IoU metric fails to effectively evaluate perception of external dynamic objects when numerous ego-vehicle points are included, we provide supplementary results on nuScenes using conventional IoU in the appendix.
+In summary, our dual-metric evaluation provides a holistic and unbiased assessment.
+The two metrics are complementary: $\text{Recall}_{\text{obj}}$ ensures that the detection completeness of each object is fairly evaluated, preventing smaller ones from being overlooked, while $\text{IoU}_{\text{w/o}}$ penalizes false positive predictions in the surrounding scene.
 5 Experiments
@@
 We perform few-shot experiments on the nuScenes dataset.
-Previous methods [34, 46] convert nuScenes object attributes into MOS labels.
+Previous methods [34, 46] directly convert nuScenes object attributes into MOS labels.
 However, these attributes do not precisely describe the object’s motion state.
-Moreover, some objects lack attributes and will be overlooked during training or evaluation.
+Moreover, some objects lack attributes and are thus overlooked during training or evaluation.
 To overcome this limitation, we refine MOS labels by incorporating object attributes and speed.
-Speed is computed from object bounding boxes and is used to classify motion states by static and moving speed thresholds.
-We introduce a speed range between these thresholds to categorize objects with unclear motion states into an unknown class, such as cars that are just starting or hesitant pedestrians.
+We compute object speed from bounding box annotations and use it to classify motion states based on predefined static and moving thresholds.
+We introduce a speed range between these thresholds to categorize objects with ambiguous motion states into an unknown class, such as cars that are just starting or hesitant pedestrians.
 More details are available in the appendix.
@@
 Implementation Details.
-In our setup, we have $n=6$, where $\mathcal{P}_{\text{adj}}$ includes 6 past scans and 6 future scans, and the input sequence $\mathcal{P}_{\text{in}}$ comprises 6 scans.
+In our setup, we set $n=6$, where $\mathcal{P}_{\text{adj}}$ includes 6 past scans and 6 future scans, and the input sequence $\mathcal{P}_{\text{in}}$ comprises 6 scans.
 All scans are sampled at 2Hz.
@@
 We apply data augmentation techniques from [37, 33] and discard points outside a bounded area based on [23, 1].
-Our encoder is the sparse 4D UNet backbone, commonly used in MOS methods [33, 34].
+Our encoder is a sparse 4D UNet.
+Following previous works [33, 34], we initialize all point features to a constant.
 The quantization size for the sparse backbone is 0.1 m, and the feature dimension $d=128$.
-For the temporal overlapping prediction objective, all occupied class samples are included due to their scarcity, while 5 times the number of occupied samples are randomly selected for the free class and the same amount for the unknown class in each training iteration.
-As for the current occupancy reconstruction objective, we select 5 occupied points and 25 free points on each current beam ($M^{\prime}=30$).
-The class weight is $\mathbf{w}_{s}=[1,5,1]$ for free, occupied, and unknown class.
+For our temporal overlapping prediction objective, we address class imbalance due to the scarcity of occupied samples.
+Specifically, in each training iteration, we use all available occupied samples, while randomly sampling five times as many points for the free class and the same number of points for the unknown class.
+As for the current occupancy reconstruction objective, we sample 5 occupied points and 25 free points on each current beam ($M^{\prime}=30$).
+The class weight is $\mathbf{w}_{s}=[1,5,1]$ for free, occupied, and unknown classes.
 We use a fixed random seed for all experiments to ensure reproducibility.
@@
 Baselines.
-We use the SOTA supervised MOS method 4DMOS [33] as the training-from-scratch baseline.
-Additionally, we implement the SOTA LiDAR observation forecasting methods 4Docc [23] and UnO [1], along with the occupancy-based 3D method ALSO [4], as the self-supervised pre-training baselines.
-All baselines utilize the same sparse 4D UNet [12, 33] encoder with their respective original decoders.
+We benchmark our method against two classes of baselines: (1) a training-from-scratch baseline using the sparse 4D UNet from [33, 34, 50], and (2) self-supervised pre-training baselines, including the LiDAR forecasting methods 4DOcc [23] and UnO [1], and the occupancy-based 3D pre-training method ALSO [4].
+To ensure a fair comparison, we replace their native encoders with the sparse 4D UNet while retaining their original decoders.
 5.2 Few-shot Results on nuScenes
-We perform few-shot MOS experiments on the nuScenes dataset.
-All methods are first pre-trained on full sequences from the nuScenes train split.
-Then the pre-trained encoder of the last epoch is chosen and fine-tuned with labeled MOS data using binary Cross-Entropy loss.
-We randomly select 5%/10%/20%/50% of sequences from the nuScenes train split for MOS fine-tuning.
-To assess the convergence performance, all models are trained for a sufficient number of epochs until convergence.
-As it is not possible to assign weights to the two metrics to select a single “best” model, we can only choose two best models separately based on each metric.
-We present the test results on full sequences from the nuScenes val split in Tab.
+We evaluate few-shot MOS on nuScenes.
+We first pre-train an encoder on the entire training split.
+Then, the encoder is fine-tuned for point-wise MOS prediction using a binary Cross-Entropy loss on randomly sampled subsets (5%, 10%, 20%, 50%) of the nuScenes training sequences.
+All models are trained to convergence.
+For evaluation, we select the best-performing checkpoint for each of our two metrics separately.
+This model selection is performed on a validation set created by randomly partitioning the official nuScenes training split.
+The final results are reported on the official nuScenes validation split (used as our test set) in Tab.
 1.
-Our pre-training method improves training-from-scratch baseline across all fine-tuning data amounts.
-Our method achieves significant improvement in $\text{mIoU}_{\text{obj}}$, up to 12.17% relative improvement from 24.98% to 28.03% $\text{mIoU}_{\text{obj}}$ when using 10% data, while the improvement in $\text{IoU}_{\text{w/o}}$ is less pronounced.
-This is because we detect more moving objects with less number of points (such as pedestrians or cycles), which account for a small proportion of the total points.
-It is noteworthy that UnO also shows some increases on $\text{mIoU}_{\text{obj}}$, while the $\text{IoU}_{\text{w/o}}$ all decreases.
-This result indicates that the performance of pre-training by UnO is not consistent, and it introduces more false negatives when detecting moving objects.
-This will lead to erroneous moving object alerts in MOS applications and cause incorrect path planning.
-In contrast, our method demonstrates consistent results in most cases.
-We visualize some results in Fig.
-4.
-Data Pretrain Best $\textbf{mIoU}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
-$\textbf{mIoU}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{mIoU}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
+TOP outperforms the training-from-scratch baseline across all data subsets.
+It demonstrates substantial gains in $\text{Recall}_{\text{obj}}$, achieving up to 12.17% relative improvement (from 24.98% to 28.03%) on the 10% data subset.
+While the gains in $\text{IoU}_{\text{w/o}}$ are more modest, we attribute this to TOP’s enhanced ability to detect small moving objects, such as pedestrians and cyclists.
+These objects significantly contribute to recall but constitute a small fraction of the total points, thus having a lesser impact on IoU.
+Data Pretrain Best $\textbf{Recall}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
+$\textbf{Recall}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{Recall}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
 5% No 24.27 34.96 23.20 35.05
@@
 4DOcc $\text{22.14}^{\text{{\color[rgb]{0.68,0,0}-2.13}}}$ $\text{33.39}^{\text{{\color[rgb]{0.68,0,0}-1.57}}}$ $\text{21.02}^{\text{{\color[rgb]{0.68,0,0}-2.17}}}$ $\text{33.81}^{\text{{\color[rgb]{0.68,0,0}-1.24}}}$
-Uno $\text{25.23}^{\text{{\color[rgb]{0,0.88,0}+0.96}}}$ $\text{34.31}^{\text{{\color[rgb]{0.68,0,0}-0.65}}}$ $\text{22.62}^{\text{{\color[rgb]{0.68,0,0}-0.58}}}$ $\text{34.90}^{\text{{\color[rgb]{0.68,0,0}-0.15}}}$
-Ours $\textbf{25.75}^{\textbf{{\color[rgb]{0,0.88,0}+1.48}}}$ $\textbf{35.94}^{\textbf{{\color[rgb]{0,0.88,0}+0.98}}}$ $\textbf{25.75}^{\textbf{{\color[rgb]{0,0.88,0}+2.56}}}$ $\textbf{35.94}^{\textbf{{\color[rgb]{0,0.88,0}+0.90}}}$
+UnO $\text{25.23}^{\text{{\color[rgb]{0,0.6,0}+0.96}}}$ $\text{34.31}^{\text{{\color[rgb]{0.68,0,0}-0.65}}}$ $\text{22.62}^{\text{{\color[rgb]{0.68,0,0}-0.58}}}$ $\text{34.90}^{\text{{\color[rgb]{0.68,0,0}-0.15}}}$
+TOP $\textbf{25.75}^{\textbf{{\color[rgb]{0,0.6,0}+1.48}}}$ $\textbf{35.94}^{\textbf{{\color[rgb]{0,0.6,0}+0.98}}}$ $\textbf{25.75}^{\textbf{{\color[rgb]{0,0.6,0}+2.56}}}$ $\textbf{35.94}^{\textbf{{\color[rgb]{0,0.6,0}+0.90}}}$
 10% No 24.98 36.44 22.98 38.06
-ALSO $\text{27.32}^{\text{{\color[rgb]{0,0.88,0}+2.34}}}$ $\text{35.74}^{\text{{\color[rgb]{0.68,0,0}-0.70}}}$ $\text{22.47}^{\text{{\color[rgb]{0.68,0,0}-0.51}}}$ $\textbf{38.33}^{\textbf{{\color[rgb]{0,0.88,0}+0.27}}}$
+ALSO $\text{27.32}^{\text{{\color[rgb]{0,0.6,0}+2.34}}}$ $\text{35.74}^{\text{{\color[rgb]{0.68,0,0}-0.70}}}$ $\text{22.47}^{\text{{\color[rgb]{0.68,0,0}-0.51}}}$ $\textbf{38.33}^{\textbf{{\color[rgb]{0,0.6,0}+0.27}}}$
 4DOcc $\text{24.31}^{\text{{\color[rgb]{0.68,0,0}-0.67}}}$ $\text{36.01}^{\text{{\color[rgb]{0.68,0,0}-0.43}}}$ $\text{23.83}^{\text{{\color[rgb]{0.68,0,0}+0.86}}}$ $\text{37.46}^{\text{{\color[rgb]{0.68,0,0}-0.70}}}$
-Uno $\text{26.38}^{\text{{\color[rgb]{0,0.88,0}+1.40}}}$ $\text{36.56}^{\text{{\color[rgb]{0,0.88,0}+0.12}}}$ $\text{22.00}^{\text{{\color[rgb]{0.68,0,0}-0.98}}}$ $\text{37.17}^{\text{{\color[rgb]{0.68,0,0}-0.89}}}$
-Ours $\textbf{28.03}^{\textbf{{\color[rgb]{0,0.88,0}+3.04}}}$ $\textbf{36.95}^{\textbf{{\color[rgb]{0,0.88,0}+0.51}}}$ $\textbf{24.69}^{\textbf{{\color[rgb]{0,0.88,0}+1.72}}}$ $\text{38.06}^{\text{{\color[rgb]{0,0.88,0}+0.00}}}$
+UnO $\text{26.38}^{\text{{\color[rgb]{0,0.6,0}+1.40}}}$ $\text{36.56}^{\text{{\color[rgb]{0,0.6,0}+0.12}}}$ $\text{22.00}^{\text{{\color[rgb]{0.68,0,0}-0.98}}}$ $\text{37.17}^{\text{{\color[rgb]{0.68,0,0}-0.89}}}$
+TOP $\textbf{28.03}^{\textbf{{\color[rgb]{0,0.6,0}+3.04}}}$ $\textbf{36.95}^{\textbf{{\color[rgb]{0,0.6,0}+0.51}}}$ $\textbf{24.69}^{\textbf{{\color[rgb]{0,0.6,0}+1.72}}}$ $\text{38.06}^{\text{{\color[rgb]{0,0.6,0}+0.00}}}$
 20% No 25.59 44.25 24.52 45.05
-ALSO $\text{26.68}^{\text{{\color[rgb]{0,0.88,0}+1.08}}}$ $\text{40.84}^{\text{{\color[rgb]{0.68,0,0}-3.41}}}$ $\text{25.81}^{\text{{\color[rgb]{0,0.88,0}+1.28}}}$ $\text{42.33}^{\text{{\color[rgb]{0.68,0,0}-2.72}}}$
-4Docc $\text{26.01}^{\text{{\color[rgb]{0,0.88,0}+0.42}}}$ $\text{42.21}^{\text{{\color[rgb]{0.68,0,0}-2.04}}}$ $\text{24.58}^{\text{{\color[rgb]{0,0.88,0}+0.06}}}$ $\text{43.11}^{\text{{\color[rgb]{0.68,0,0}-1.94}}}$
-Uno $\text{28.14}^{\text{{\color[rgb]{0,0.88,0}+2.55}}}$ $\text{43.63}^{\text{{\color[rgb]{0.68,0,0}-0.62}}}$ $\text{26.78}^{\text{{\color[rgb]{0,0.88,0}+2.26}}}$ $\text{43.97}^{\text{{\color[rgb]{0.68,0,0}-1.08}}}$
-Ours $\textbf{28.45}^{\textbf{{\color[rgb]{0,0.88,0}+2.86}}}$ $\textbf{44.30}^{\textbf{{\color[rgb]{0,0.88,0}+0.05}}}$ $\textbf{27.29}^{\textbf{{\color[rgb]{0,0.88,0}+2.76}}}$ $\text{44.52}^{\text{{\color[rgb]{0.68,0,0}-0.53}}}$
+ALSO $\text{26.68}^{\text{{\color[rgb]{0,0.6,0}+1.08}}}$ $\text{40.84}^{\text{{\color[rgb]{0.68,0,0}-3.41}}}$ $\text{25.81}^{\text{{\color[rgb]{0,0.6,0}+1.28}}}$ $\text{42.33}^{\text{{\color[rgb]{0.68,0,0}-2.72}}}$
+4DOcc $\text{26.01}^{\text{{\color[rgb]{0,0.6,0}+0.42}}}$ $\text{42.21}^{\text{{\color[rgb]{0.68,0,0}-2.04}}}$ $\text{24.58}^{\text{{\color[rgb]{0,0.6,0}+0.06}}}$ $\text{43.11}^{\text{{\color[rgb]{0.68,0,0}-1.94}}}$
+UnO $\text{28.14}^{\text{{\color[rgb]{0,0.6,0}+2.55}}}$ $\text{43.63}^{\text{{\color[rgb]{0.68,0,0}-0.62}}}$ $\text{26.78}^{\text{{\color[rgb]{0,0.6,0}+2.26}}}$ $\text{43.97}^{\text{{\color[rgb]{0.68,0,0}-1.08}}}$
+TOP $\textbf{28.45}^{\textbf{{\color[rgb]{0,0.6,0}+2.86}}}$ $\textbf{44.30}^{\textbf{{\color[rgb]{0,0.6,0}+0.05}}}$ $\textbf{27.29}^{\textbf{{\color[rgb]{0,0.6,0}+2.76}}}$ $\text{44.52}^{\text{{\color[rgb]{0.68,0,0}-0.53}}}$
 50% No 29.80 49.29 29.80 49.29
-ALSO $\text{30.72}^{\text{{\color[rgb]{0,0.88,0}+0.91}}}$ $\text{47.34}^{\text{{\color[rgb]{0.68,0,0}-1.95}}}$ $\text{28.06}^{\text{{\color[rgb]{0.68,0,0}-1.75}}}$ $\text{48.10}^{\text{{\color[rgb]{0.68,0,0}-1.19}}}$
+ALSO $\text{30.72}^{\text{{\color[rgb]{0,0.6,0}+0.91}}}$ $\text{47.34}^{\text{{\color[rgb]{0.68,0,0}-1.95}}}$ $\text{28.06}^{\text{{\color[rgb]{0.68,0,0}-1.75}}}$ $\text{48.10}^{\text{{\color[rgb]{0.68,0,0}-1.19}}}$
 4DOcc $\text{29.20}^{\text{{\color[rgb]{0.68,0,0}-0.60}}}$ $\text{46.73}^{\text{{\color[rgb]{0.68,0,0}-2.56}}}$ $\text{28.19}^{\text{{\color[rgb]{0.68,0,0}-1.62}}}$ $\text{48.19}^{\text{{\color[rgb]{0.68,0,0}-1.10}}}$
 UnO $\text{29.54}^{\text{{\color[rgb]{0.68,0,0}-0.26}}}$ $\text{48.59}^{\text{{\color[rgb]{0.68,0,0}-0.70}}}$ $\text{28.81}^{\text{{\color[rgb]{0.68,0,0}-1.00}}}$ $\text{48.76}^{\text{{\color[rgb]{0.68,0,0}-0.53}}}$
-Ours $\textbf{30.58}^{\textbf{{\color[rgb]{0,0.88,0}+0.78}}}$ $\textbf{49.80}^{\textbf{{\color[rgb]{0,0.88,0}+0.51}}}$ $\textbf{30.58}^{\textbf{{\color[rgb]{0,0.88,0}+0.78}}}$ $\textbf{49.80}^{\textbf{{\color[rgb]{0,0.88,0}+0.51}}}$
-Table 1: Few-shot MOS results on the nuScenes dataset. The best performance in each column is in bold. The metric used for model selection is in bold in the table header.
-5.3 Transferring Results on SemanticKITTI
-To evaluate the transferability of the learned spatiotemporal representations, we fine-tune the pre-trained 4D UNet encoder on the SemanticKITTI dataset.
-We train the model on 10%/20%50% random samples from SemanticKITTI train set for enough epochs until convergence.
-As demonstrated in Tab.
-2, TOP shows the best transferability on different LiDAR setups among all baselines.
-Fig.
-5 displays the results on the SemanticKITTI dataset.
-Data Pretrain Best $\textbf{mIoU}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
-$\textbf{mIoU}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{mIoU}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
+TOP $\textbf{30.58}^{\textbf{{\color[rgb]{0,0.6,0}+0.78}}}$ $\textbf{49.80}^{\textbf{{\color[rgb]{0,0.6,0}+0.51}}}$ $\textbf{30.58}^{\textbf{{\color[rgb]{0,0.6,0}+0.78}}}$ $\textbf{49.80}^{\textbf{{\color[rgb]{0,0.6,0}+0.51}}}$
+Table 1: Few-shot MOS results on nuScenes. The best performance in each column is in bold. The metric used for model selection is in bold in the table header.
+5.3 Few-shot Transfer Results on SemanticKITTI
+We evaluate the transferability of our learned spatiotemporal representations by fine-tuning the pre-trained encoder on SemanticKITTI in a few-shot setting.
+Specifically, we use random subsets of 10%, 20%, and 50% of the official training set to fine-tune the model until convergence.
+We report the final results on the official SemanticKITTI validation set.
+As shown in Tab.
+2, TOP demonstrates superior transferability, significantly outperforming the training-from-scratch baseline and other pre-training baselines across all data subsets.
+We present qualitative MOS results on nuScenes and SemanticKITTI in Fig.
+4 and Fig.
+5, respectively.
+Data Pretrain Best $\textbf{Recall}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
+$\textbf{Recall}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{Recall}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
 10% No 37.22 41.22 33.97 42.81
 ALSO $\text{9.21}^{\text{{\color[rgb]{0.68,0,0}-28.01}}}$ $\text{7.33}^{\text{{\color[rgb]{0.68,0,0}-33.88}}}$ $\text{9.21}^{\text{{\color[rgb]{0.68,0,0}-24.77}}}$ $\text{7.76}^{\text{{\color[rgb]{0.68,0,0}-35.05}}}$
-4Docc $-28.68\text{8.54}^{\text{{\color[rgb]{0.68,0,0}{-28.68}}}}$ $-27.87\text{13.35}^{\text{{\color[rgb]{0.68,0,0}{-27.87}}}}$ $-25.82\text{8.15}^{\text{{\color[rgb]{0.68,0,0}{-25.82}}}}$ $\text{13.45}^{\text{{\color[rgb]{0.68,0,0}-29.36}}}$
-Uno $-0.57\text{36.66}^{\text{{\color[rgb]{0.68,0,0}{-0.57}}}}$ $-6.19\text{35.03}^{\text{{\color[rgb]{0.68,0,0}{-6.19}}}}$ $-6.15\text{27.83}^{\text{{\color[rgb]{0.68,0,0}{-6.15}}}}$ $\text{37.41}^{\text{{\color[rgb]{0.68,0,0}-5.41}}}$
-Ours $\textbf{44.68}^{\textbf{{\color[rgb]{0,0.88,0}+7.46}}}$ $\textbf{43.29}^{\textbf{{\color[rgb]{0,0.88,0}+2.07}}}$ $\textbf{43.36}^{\textbf{{\color[rgb]{0,0.88,0}+9.39}}}$ $\textbf{49.37}^{\textbf{{\color[rgb]{0,0.88,0}+6.55}}}$
+[…]
+[…]
+TOP $\textbf{44.68}^{\textbf{{\color[rgb]{0,0.6,0}+7.46}}}$ $\textbf{43.29}^{\textbf{{\color[rgb]{0,0.6,0}+2.07}}}$ $\textbf{43.36}^{\textbf{{\color[rgb]{0,0.6,0}+9.39}}}$ $\textbf{49.37}^{\textbf{{\color[rgb]{0,0.6,0}+6.55}}}$
 20% No 47.41 47.38 45.39 48.77
 ALSO $\text{23.37}^{\text{{\color[rgb]{0.68,0,0}-24.04}}}$ $\text{23.02}^{\text{{\color[rgb]{0.68,0,0}-24.36}}}$ $\text{23.37}^{\text{{\color[rgb]{0.68,0,0}-22.02}}}$ $\text{23.02}^{\text{{\color[rgb]{0.68,0,0}-25.75}}}$
-4Docc $-21.80\text{25.61}^{\text{{\color[rgb]{0.68,0,0}{-21.80}}}}$ $-11.55\text{35.83}^{\text{{\color[rgb]{0.68,0,0}{-11.55}}}}$ $-20.16\text{25.23}^{\text{{\color[rgb]{0.68,0,0}{-20.16}}}}$ $\text{36.24}^{\text{{\color[rgb]{0.68,0,0}-12.53}}}$
+[…]
 UnO $\text{45.48}^{\text{{\color[rgb]{0.68,0,0}-1.92}}}$ $\text{42.09}^{\text{{\color[rgb]{0.68,0,0}-5.29}}}$ $\text{43.77}^{\text{{\color[rgb]{0.68,0,0}-1.62}}}$ $\text{46.09}^{\text{{\color[rgb]{0.68,0,0}-2.68}}}$
-Ours $\textbf{58.45}^{\textbf{{\color[rgb]{0,0.88,0}+11.04}}}$ $\textbf{50.21}^{\textbf{{\color[rgb]{0,0.88,0}+2.82}}}$ $\textbf{58.45}^{\textbf{{\color[rgb]{0,0.88,0}+13.06}}}$ $\textbf{50.21}^{\textbf{{\color[rgb]{0,0.88,0}+1.44}}}$
+TOP $\textbf{58.45}^{\textbf{{\color[rgb]{0,0.6,0}+11.04}}}$ $\textbf{50.21}^{\textbf{{\color[rgb]{0,0.6,0}+2.82}}}$ $\textbf{58.45}^{\textbf{{\color[rgb]{0,0.6,0}+13.06}}}$ $\textbf{50.21}^{\textbf{{\color[rgb]{0,0.6,0}+1.44}}}$
 50% No 53.13 56.35 48.20 59.20
-ALSO $\text{49.82}^{\text{{\color[rgb]{0.68,0,0}-3.32}}}$ $\text{57.07}^{\text{{\color[rgb]{0,0.88,0}+0.72}}}$ $\text{44.27}^{\text{{\color[rgb]{0.68,0,0}-3.93}}}$ $\text{60.25}^{\text{{\color[rgb]{0,0.88,0}+1.04}}}$
+ALSO $\text{49.82}^{\text{{\color[rgb]{0.68,0,0}-3.32}}}$ $\text{57.07}^{\text{{\color[rgb]{0,0.6,0}+0.72}}}$ $\text{44.27}^{\text{{\color[rgb]{0.68,0,0}-3.93}}}$ $\text{60.25}^{\text{{\color[rgb]{0,0.6,0}+1.04}}}$
 4DOcc $\text{39.23}^{\text{{\color[rgb]{0.68,0,0}-13.91}}}$ $\text{44.48}^{\text{{\color[rgb]{0.68,0,0}-11.88}}}$ $\text{38.36}^{\text{{\color[rgb]{0.68,0,0}-9.84}}}$ $\text{44.87}^{\text{{\color[rgb]{0.68,0,0}-14.33}}}$
-Uno $\text{53.61}^{\text{{\color[rgb]{0,0.88,0}+0.48}}}$ $\text{54.97}^{\text{{\color[rgb]{0.68,0,0}-1.38}}}$ $\textbf{52.73}^{\textbf{{\color[rgb]{0,0.88,0}+4.53}}}$ $\text{57.06}^{\text{{\color[rgb]{0.68,0,0}-2.14}}}$
-Ours $\textbf{55.81}^{\textbf{{\color[rgb]{0,0.88,0}+2.67}}}$ $\textbf{58.85}^{\textbf{{\color[rgb]{0,0.88,0}+2.50}}}$ $\text{52.11}^{\text{{\color[rgb]{0,0.88,0}+3.91}}}$ $\textbf{61.55}^{\textbf{{\color[rgb]{0,0.88,0}+2.35}}}$
-Table 2: Cross-dataset transferring results on the SemanticKITTI MOS benchmark. The best performance in each column is in bold. The metric used for model selection is in bold in the table header.
+UnO $\text{53.61}^{\text{{\color[rgb]{0,0.6,0}+0.48}}}$ $\text{54.97}^{\text{{\color[rgb]{0.68,0,0}-1.38}}}$ $\textbf{52.73}^{\textbf{{\color[rgb]{0,0.6,0}+4.53}}}$ $\text{57.06}^{\text{{\color[rgb]{0.68,0,0}-2.14}}}$
+TOP $\textbf{55.81}^{\textbf{{\color[rgb]{0,0.6,0}+2.67}}}$ $\textbf{58.85}^{\textbf{{\color[rgb]{0,0.6,0}+2.50}}}$ $\text{52.11}^{\text{{\color[rgb]{0,0.6,0}+3.91}}}$ $\textbf{61.55}^{\textbf{{\color[rgb]{0,0.6,0}+2.35}}}$
+Table 2: Cross-dataset few-shot transfer results on SemanticKITTI. The best performance in each column is in bold. The metric used for model selection is in bold in the table header.
+Figure 4: Qualitative results of the nuScenes MOS.
+Figure 5: Qualitative results of the SemanticKITTI MOS.
 5.4 Semantic Segmentation Results on nuScenes
-We also assess how well our pre-training generalizes to a new downstream task by conducting a few-shot semantic segmentation experiment on the nuScenes dataset.
-We fine-tune the pre-trained encoders for a sufficient number of epochs using 10% of the train split sequences.
+We evaluate the generalization of our pre-training on a few-shot semantic segmentation task on nuScenes, by fine-tuning the pre-trained encoder using 10% of the training split.
 As shown in Tab.
-3, TOP shows a +0.42% mIoU improvement compared to the training-from-scratch baseline, while the performance of other methods notably declines.
-We suspect that the performance drop of 4Docc and UnO may be due to the complex “flow” information causing network overfitting, as discussed in Sec.
-1.
-This experiment demonstrates the potential generalization capability of TOP pre-training across different downstream tasks.
-Pretrain No ALSO 4Docc UnO Ours
-mIoU 50.55 $50.26^{\textbf{{\color[rgb]{0.68,0,0}-0.29}}}$ $47.49^{\textbf{{\color[rgb]{0.68,0,0}-3.06}}}$ $48.06^{\textbf{{\color[rgb]{0.68,0,0}-2.49}}}$ $\textbf{50.97}^{\textbf{{\color[rgb]{0,0.88,0}+0.42}}}$
+3, TOP surpasses the training-from-scratch baseline by 0.42% mIoU at convergence.
+In contrast, models pre-trained with 4DOcc and UnO exhibit a notable performance drop.
+We attribute this drop to their overfitting to the complex “flow” information during their pre-training (Sec.
+1).
+This result demonstrates the generalization potential of TOP pre-training across different downstream tasks.
+Pretrain No ALSO 4DOcc UnO TOP
+mIoU 50.55 $50.26^{\text{{\color[rgb]{0.68,0,0}-0.29}}}$ $47.49^{\text{{\color[rgb]{0.68,0,0}-3.06}}}$ $48.06^{\text{{\color[rgb]{0.68,0,0}-2.49}}}$ $\textbf{50.97}^{\textbf{{\color[rgb]{0,0.6,0}+0.42}}}$
 Table 3: Few-shot results on nuScenes semantic segmentation.
@@
 As current occupancy reconstruction is straightforward and its loss converges rapidly, we speculate it acts as a regularization term to prevent overfitting.
-Objectives Best $\textbf{mIoU}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
-$\textbf{mIoU}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{mIoU}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
-$\mathcal{L}_{\text{overlap}}$ 24.559 34.785 22.639 35.350
+Objectives Best $\textbf{Recall}_{\textbf{obj}}$ Best $\textbf{IoU}_{\textbf{w/o}}$
+$\textbf{Recall}_{\textbf{obj}}$ $\text{IoU}_{\text{w/o}}$ $\text{Recall}_{\text{obj}}$ $\textbf{IoU}_{\textbf{w/o}}$
+$\mathcal{L}_{\text{overlap}}$ 24.56 34.79 22.64 35.35
 $\mathcal{L}_{\text{overlap}}+\mathcal{L}_{\text{recon}}$ 25.75 35.94 25.75 35.94
-Table 4: Ablation study on two training objectives.
-Figure 4: Visualization of the nuScenes MOS results.
-Figure 5: Visualization of the SemanticKITTI MOS results.
+Table 4: Ablation study on training objectives.
 6 Conclusion
-In this paper, we propose a self-supervised pre-training method called Temporal Overlapping Prediction (TOP) for moving object segmentation (MOS).
-TOP predicts the occupancy states of temporal overlapping points to help the model learn spatiotemporal representations for MOS.
-We evaluate the effectiveness of TOP through few-shot experiments on nuScenes and cross-dataset transfer experiments on SemanticKITTI.
-TOP consistently improves both the object-level metric and the point-level metric, demonstrating strong transferability across LiDAR setups.
-Additionally, the improved performance in semantic segmentation suggests potential generalization to more downstream tasks.
-Further investigation of the effectiveness of TOP in other temporal tasks such as object tracking is worthwhile.
+In this paper, we propose Temporal Overlapping Prediction (TOP), a self-supervised pre-training method for MOS.
+By learning to predict the occupancy of temporal overlapping points, TOP develops powerful spatiotemporal representations for moving object segmentation.
+We evaluate its effectiveness through few-shot experiments on nuScenes and cross-dataset transfer experiments on SemanticKITTI.
+TOP consistently improves both object-level and point-level metrics, demonstrating strong transferability across LiDAR setups.
+Additionally, the improved performance in semantic segmentation suggests its potential to generalize to other downstream tasks.
+We believe investigating TOP for other temporal tasks, such as object tracking, is a promising avenue for future work.
+Acknowledgement
+The computations were performed using research computing facilities offered by Information Technology Services, the University of Hong Kong.
+Supplementary Material
+In this appendix, we describe a complementary formulation, the MOS labeling process, the cumulative distribution of moving objects and points, implementation details, the impact of ego-vehicle points on IoU, and potential negative social impacts.
+Finally, we present visual examples of the MOS results on the nuScenes and SemanticKITTI datasets in Fig.
+7 and Fig.
+8, respectively.
+1 Formulation
+As described in Sec.
+3.2 Scenario-2 in the main paper, we provide the formula for the starting point $\mathbf{b}_{i}^{t_{0}}$ of the intersection segment.
+Since the spatial angle $\alpha_{i,j}$ is sufficiently small, we approximate that: (1) The two intersection segments are assumed to be equal in length, which is described as:
+[…]
+(2) The distance between the two starting points $\mathbf{b}_{i}^{t_{0}}$ and $\mathbf{b}_{j}^{t}$ is equal to the sum of beam radii at the two points, represented as follows:
+[…]
+We substitute Eq. 13 into Eq. 14 and solve for $\mathbf{b}_{i}^{t_{0}}||$:
+[…]
+2 MOS Labeling
+As discussed in Sec.
+5.1 of the main paper, the nuScenes object attributes do not precisely describe the motion state, for example, “cycle.with_rider” can be either moving or static.
+Furthermore, a proportion of objects lack attributes: $0.466\%$ of vehicles, $9.243\%$ of cycles, and $1.913\%$ of pedestrians.
+For our nuScenes MOS labeling, we calculate object speeds from their bounding box annotations.
+An object is classified as static if its speed is less than $\mu_{\text{sta}}$, and as moving if its speed exceeds $\mu_{\text{mov}}$.
+Objects with speeds between these thresholds are considered to have an unclear motion state and are thus classified as unknown.
+We use different speed thresholds for humans $(\mu_{\text{sta}}^{\text{hum}}=0.375,\mu_{\text{mov}}^{\text{hum}}=0.6)$, cycles $(\mu_{\text{sta}}^{\text{cyc}}=0.375,\mu_{\text{mov}}^{\text{cyc}}=1.0)$, and vehicles $(\mu_{\text{sta}}^{\text{veh}}=0.5,\mu_{\text{mov}}^{\text{veh}}=1.0)$.
+3 Cumulative Distribution
+Fig.
+6 presents a statistical analysis of moving objects from the nuScenes train-val split (discussed in Sec.
+4), plotting the cumulative distribution functions of moving object instances and the number of their scanned points.
+The distributions are analyzed with respect to object size, defined as the number of points per object (x-axis).
+The significant gap between the two curves highlights a strong imbalance: the vast majority of moving objects consist of very few points.
+This is evidenced by the data: while moving objects with 19 or fewer points comprise 75.22% of all moving objects, they contribute a mere 15.49% of the total moving points.
+This disparity grows, with the smallest 90.06% of objects collectively accounting for only 37.02% of the total points.
+Such a skewed distribution implies that a small number of point-rich objects can disproportionately dominate the conventional IoU metric.
+Figure 6: Cumulative distribution of moving objects and points.
+4 Implementation Details
+method epochs batch size optimizer lr start lr max
+TOP 50 8 AdamW $5.0\times 10^{-6}$ $5.0\times 10^{-5}$
+ALSO 200 4 AdamW $\times 10^{-4} 1.0×10$$\times 10^{-3}$
+4DOcc 50 2 AdamW $8.0\times 10^{-5}$ $8.0\times 10^{-4}$
+UnO 50 16 AdamW $8.0\times 10^{-5}$ $8.0\times 10^{-4}$
+Pre-training Settings.
+The detailed pre-training settings are shown in the table above.
+The learning rate (lr) follows a warmup-cosine schedule, increasing from “lr start” to “lr max” during the first 2% of iterations (the warmup stage), and then follows a cosine schedule for the remaining iterations.
+All pre-training baseline methods maintain their original settings, while 4DOcc uses a very small batch size due to the high GPU memory consumption caused by the 0.1 m dense grid map.
+Fine-tuning Settings.
+The Adam optimizer is used for all the fine-tuning; the learning rate follows a step-decay schedule, starting from $1.0\times 10^{-4}$ and decreasing by a factor of $0.99$ after each iteration.
+For the nuScenes MOS experiments, the batch size is 4 for 5%, 10%, and 20% data subsets, while the batch size for the 50% data subset is 8.
+All methods are trained for over 400 epochs to ensure convergence.
+For SemanticKITTI cross-dataset transfer experiments, the batch size is 2 for all data subsets, and all methods are trained for over 300 epochs.
+For the nuScenes semantic segmentation experiment, all methods are trained for 300 epochs with a batch size of 4.
+5 Impact of Ego-Vehicle Points on IoU Metric
+As discussed in Sec.
+4 of the main paper, the conventional IoU metric (denoted as IoU hereafter) includes ego-vehicle points, which inflates scores and masks actual perception performance on external moving objects.
+This issue is especially pronounced in the nuScenes dataset due to its high proportion of ego-vehicle points.
+To demonstrate this effect, we present nuScenes MOS results using IoU in Tab.
+5.
+The results show that IoU scores are inflated by 1.5-2x compared to the ego-exclusive metric $\text{IoU}_{\text{w/o}}$.
+Furthermore, the apparent IoU performance gains are deceptive.
+These improvements are not from enhanced perception of the external environment, but from the far simpler task of ego-vehicle segmentation.
+Therefore, relying on this metric risks severely misrepresenting a model’s actual capability.
+Data Pretrain Best $\textbf{Recall}_{\textbf{obj}}$
+$\text{Recall}_{\text{obj}}$ $\text{IoU}_{\text{w/o}}$ IoU
+10% No 24.98 36.44 68.89
+TOP $\text{28.03}^{\text{{\color[rgb]{0,0.6,0}+3.04}}}$ $\text{36.95}^{\text{{\color[rgb]{0,0.6,0}+0.51}}}$ $\text{70.28}^{\text{{\color[rgb]{0,0.6,0}+1.39}}}$
+20% No 25.59 44.25 63.68
+TOP $\text{28.45}^{\text{{\color[rgb]{0,0.6,0}+2.86}}}$ $\text{44.30}^{\text{{\color[rgb]{0,0.6,0}+0.05}}}$ $\text{68.13}^{\text{{\color[rgb]{0,0.6,0}+4.45}}}$
+Table 5: nuScenes MOS results using the conventional IoU metric.
+6 Potential Negative Social Impacts
+While the proposed self-supervised method for MOS reduces annotation costs for autonomous systems, potential societal impacts should be considered.
+The technology could unintentionally prioritize detection accuracy for dominant object classes (e.g., vehicles) over vulnerable road users like cyclists/pedestrians if trained on imbalanced datasets, potentially compromising safety in edge cases.
+The temporal nature of our approach might propagate motion prediction errors in complex urban scenarios, leading to hazardous decisions by autonomous vehicles.
+Furthermore, while addressing sensor bias through our new metric helps, residual geographic/cultural biases in training data (e.g., urban vs rural environments) could limit global applicability.
+The method could also be repurposed for surveillance systems that infringe on privacy.
+To mitigate these risks, we recommend: (1) Rigorous testing across diverse operational domains.
+(2) Implementing fairness-aware data sampling strategies.
+(3) Establishing ethical guidelines for secondary applications.
+Figure 7: Qualitative results of the nuScenes MOS.
+Figure 8: Qualitative results of the SemanticKITTI MOS.